### v1

Abstract
We develop a decoupled, first-order, fully discrete, energy-stable scheme for the Cahn-Hilliard-Navier-Stokes equations. This scheme calculates the Cahn-Hilliard and Navier-Stokes equations separately, thus effectively decoupling the entire system. To further separate the velocity and pressure components in the Navier-Stokes equations, we use the pressure-correction projection method. We demonstrate that the scheme is primitively energy stable and prove the optimal $L^{2}$ error estimate of the fully discrete scheme in the $P_{r}\times P_{r}\times P_{r}\times P_{r-1}$ finite element spaces, where the phase field, chemical potential, velocity and pressure satisfy the first-order accuracy in time and the $\left(r+1,r+1,r+1,r\right)th$-order accuracy in space, respectively. Furthermore, numerical experiments are conducted to support these theoretical findings. Notably, compared to other numerical schemes, our algorithm is more time-efficient and numerically shown to be unconditionally stable.
Keywords: Phase field models; Cahn-Hilliard-Navier-Stokes; convex-splitting; energy stability; optimal error estimates.
1 Introduction
The Cahn-Hilliard-Navier-Stokes (CHNS) model features a nonlinear interaction between the incompressible Navier-Stokes (NS) equations [1] and the Cahn-Hilliard (CH) equations [2]. This model captures the interfacial dynamics of two-phase, incompressible, and macroscopically immiscible Newtonian fluids that exhibit matched densities [3, 4] as follows:
$\frac{\partial\phi}{\partial t}+(\mathbf{u}\cdot\nabla)\phi-M\Delta\mu=0,\quad\text{in}\quad\Omega\times(0,T],$ (1.1a)
$\mu+\lambda\Delta\phi-\lambda F^{\prime}(\phi)=0,\quad\text{in}\quad\Omega\times(0,T],$ (1.1b)
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}-\nu\Delta\mathbf{u}+\nabla p-\mu\nabla\phi=0,\quad\text{in}\quad\Omega\times(0,T],$ (1.1c)
$\nabla\cdot\mathbf{u}=0,\quad\text{in}\quad\Omega\times(0,T],$ (1.1d)
with the following boundary and initial conditions of (1.1):
$\frac{\partial\phi}{\partial\mathbf{n}}=\frac{\partial\mu}{\partial\mathbf{n}}=0,\quad\mathbf{u}=0,$ $\quad\text{on}\quad\partial\Omega\times(0,T],$ (1.2a)
$\phi(\mathbf{x},0)=\phi^{0},\quad\mathbf{u}(\mathbf{x},0)=\mathbf{u}^{0},$ $\quad\text{in}\quad\Omega.$ (1.2b)
Here $\Omega\subset\mathbb{R}^{d},~{}(d=2,3)$ is a bounded convex polygonal or polyhedral domain, and for $t\in(0,T]$, and $F(\phi)=\frac{1}{4\epsilon^{2}}\left(\phi^{2}-1\right)^{2}$. In the model, $\phi$ represents the phase field variable, and $\mu$ denotes the chemical potential, and the $\mathbf{u}$ and $p$ are the velocity and pressure of the fluid, respectively. $\epsilon$ is the interfacial width between the two phases field, and the parameters $M$, $\lambda$ and $\nu$ represent the mobility constant, the mixing coefficient and the fluid viscosity, respectively. It is well know that the energy of CHNS model is defined by
[…]
The system (1.1)-(1.2) satisfies the following energy dissipation law at $\forall t\in(0,T]$,
$\frac{\partial E(\phi(t),\mathbf{u}(t))}{\partial t}=-M\|\nabla\mu\|^{2}-\nu\|\nabla\mathbf{u}\|^{2}.$ (1.4)
Over the past decade or so, one of the primary challenges for the CH equations have been the efficient treatment of nonlinear terms to ensure that the resulting discretized system is solved efficiently while maintaining energy stability. Some of the techniques for dealing with the nonlinear terms in the numerical format of the CH equation include the convex splitting [5], stabilized semi-implicit methods [6], invariant energy quadratization (IEQ) [7], and the scalar auxiliary variable (SAV) approach [8, 9]. On the other hand, the main challenges in solving the NS equations lie in the treatment of nonlinear terms and the coupling between velocity and pressure under incompressible conditions. For the coupling between velocity and pressure, a widely adopted strategy is the projection method [10], originally proposed by Chorin and Temam [11]. The nonlinear terms in the NS equations can be handled either in an fully implicit, semi-implicit, or through SAV full explicatization. Wang et al. [12] presented a novel second order in time mixed finite element scheme for the Cahn–Hilliard–Navier–Stokes equations with matched densities, and they [13] proposed and analyze a time second-order accurate numerical scheme for the Cahn–Hilliard-Magnetohydrodynamics equations. As a result, based on the several challenges mentioned earlier, a large number of scholars have worked on proposing energy-stable and efficient numerical schemes for CHNS [14, 15, 16, 17, 18, 19, 20, 21, 22, 12, 23].
Recently, Xu et al. [20] proposed a first-order in time, linear, fully decoupled, and energy stable scheme for the CHNS model. They introduced a velocity intermediate quantity in the discretization of the CH equation, thus explicitly dealing with the coupling term between the velocity and phase functions. The introduction of this intermediate variable makes error estimation involved, especially if one also wants to obtain optimal $L^{2}$ error estimates for the spatial variables. Then, Wen et al. [24] proposed a novel fully discrete stabilized finite element method using the lowest equal-order finite elements, and they analyzed the optimal $H^{1}$ error estimates for the phase field function, chemical potential and velocity. Cai et al. [25] first presented optimal $L^{2}$ error estimates of the convex splitting finite element method for the CHNS system. However, these schemes are a coupled system, whose numerical results in relatively long computation time. Soon after, Yang et al. [26] proposed a step-by-step decoupling method that utilizes the SAV approach to split the CHNS model into CH and NS components, enabling the solution of these two equations separately, and also analyzed the optimal $L^{2}$ error estimates for the numerical solution, but their scheme did not achieve original energy stabilization.
To the best of our knowledge, there is no literature on the optimal $L^{2}$ error estimates for the fully decoupled, original energy stable numerical scheme of CHNS model. To achieve this purpose, we firstly decouple the CH and NS equations by handling the convective term implicitly/explicitly in the CH equation, thus resulting in a fully decoupled and original energy stable semi-discrete scheme. Secondly, we discretize the spatial variables using the mixed finite element method to obtain a fully decoupled, fully discretized scheme with original energy stable, which is analyzed to obtain the optimal $L^{2}$ error estimate. We notice that Kay et al. [4] also constructed a fully decoupled numerical scheme with original energy stabilization, they, however, did not provide the optimal $L^{2}$ error estimates. Additionally, Cai et al. [25] analytically derived the optimal $L^{2}$ error estimates for the scheme with original energy stabilization, but the coupled nature of their scheme made computation expensive. We emphasize that despite a mild time-step restriction is imposed in our energy stability, our numerical experiment demonstrates that our fully decoupled scheme is unconditionally energy stable.
This manuscript is organized as follows: The Section 2 introduces fundamental notations and lemmas. In Section 3, we propose a fully decoupled, energy-stable scheme utilizing fully discrete mixed finite element methods based on pressure-correction and convex splitting techniques for the CHNS model; we then prove the energy stability of our scheme. In Section 4, we provide the corresponding error estimates and establish optimal $L^{2}$ error estimates for the fully discrete scheme. Numerical experiments are conducted in Section 5 to validate the theoretical results of our scheme.
2 The mathematical prelimilaries
[…]
$b(\phi,\bm{v},w)=\left(\nabla\phi\cdot\bm{v},w\right),$ (2.1)
and for $\bm{u},\bm{v},\bm{w}\in\mathbf{H}^{1}(\Omega)$, we define the trilinear form $\bm{B}(\cdot,\cdot,\cdot)$ by
$\bm{B}(\bm{u},\bm{v},\bm{w})=\frac{1}{2}\left(\left(\bm{u}\cdot\nabla\right)\bm{v},\bm{w}\right)-\frac{1}{2}\left((\bm{u}\cdot\nabla)\bm{w},\bm{v}\right).$ (2.2)
It is well-know result that
$\bigg{|}b(\phi,\bm{v},w)\bigg{|}=\bigg{|}\int_{\Omega}w\nabla\phi\cdot\bm{v}dx\bigg{|}\leq\|\phi\bm{v}\|\|\nabla w\|\leq\|\phi\|_{L^{4}}\|\bm{v}\|_{L^{4}}\|\nabla w\|\leq\|\phi\|_{H^{1}}\|\bm{v}\|_{H^{1}}\|\nabla w\|.$ (2.3)
We next introduce the Ritz and Stokes projections. Let $\mathcal{T}_{h}$ denotes a quasi-uniform partition of $\Omega$ into triangles $\mathcal{K}_{j},~{}j=1,\cdots,M$, in $\mathbb{R}^{2}$ with mesh size $h=\max_{1\leq j\leq M}\{\dim\mathcal{K}_{j}\}$. For a given positive integer $r\geq 2$, we set the phase field-velocity-pressure finite element spaces
$S_{h}^{r}=\left\{v_{h}\in C(\Omega):\left.v_{h}\right|_{\mathcal{K}_{j}}\in P_{r}\left(\mathcal{K}_{j}\right),\forall\mathcal{K}_{j}\in\mathcal{T}_{h}\right\},$ (2.4)
$\mathring{S}_{h}^{r-1}=S_{h}^{r-1}\cap L_{0}^{2}(\Omega),$
$\mathbf{X}_{h}^{r}=\left\{\mathbf{v}_{h}\in\mathbf{H}_{0}^{1}(\Omega)^{d}:\left.\mathbf{v}_{h}\right|_{\mathcal{K}_{j}}\in\mathbf{P}_{r}\left(\mathcal{K}_{j}\right)^{d},\forall\mathcal{K}_{j}\in\mathcal{T}_{h}\right\},$
where $P_{r}\left(\mathcal{K}_{j}\right)$ is the space of polynomials of degree $r$ on $\mathcal{K}_{j}$. For some constant $C>0$, it is well-known that both the Taylor-Hood elements satisfy the discrete inf-sup condition [27]
$\left\|q_{h}\right\|_{L^{2}}\leq C\sup_{\mathbf{0}\neq\mathbf{v}_{h}\in\mathbf{X}_{h}^{r}}\frac{\left(q_{h},\nabla\cdot\mathbf{v}_{h}\right)}{\left\|\nabla\mathbf{v}_{h}\right\|},\quad\forall q_{h}\in\mathring{S}_{h}^{r-1}.$ (2.5)
For the simplicity of notations and $r\geq 2$, we define
$\mathbf{\mathcal{X}}_{h}^{r}=S_{h}^{r}\times S_{h}^{r}\times\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1},$ (2.6)
and denote $\bm{g}^{n+1}=\bm{g}\left(x,t^{n+1}\right)$, and
$\delta_{\tau}\bm{g}^{n+1}=\frac{\bm{g}^{n+1}-\bm{g}^{n}}{\tau}.$ (2.7)
Remark 2.1.
In an abuse of notation, we use $\bm{g}^{n+1}$ hereafter to denote the value of the exact solution $\bm{g}$ at $t^{n+1}$.
We will frequently use the following discrete version of the Grönwall lemma [28]:
Lemma 2.1 (discrete Grönwall inequality [28]).
For all $0\leq n\leq m$, let $a_{n},b_{n},c_{n},d_{n},\tau,C\geq 0$ such that
$a_{m}+\tau\sum_{n=0}^{m}b_{n}\leq\tau\sum_{n=0}^{m}d_{n}a_{n}+\tau\sum_{n=0}^{m}c_{n}+C,$ (2.8)
suppose that $\tau d_{n}<1$, for all $n\geq 0$, and set $\sigma_{n}=(1-\tau d_{n})^{-1}$. Then
$a_{m}+\tau\sum_{n=0}^{m}b_{n}\leq\exp\left(\tau\sum_{n=0}^{m}\sigma_{n}d_{n}\right)\left\{a_{0}+\left(b_{0}+c_{0}\right)\tau+\tau\sum_{n=1}^{m}c_{n}\right\}.$ (2.9)
Throughout the manuscript we use $C$ or $c$, with or without subscript, to denote a positive constant independent of discretization parameters that could have uncertain values in different places.
3 The numerical scheme
In the section, we construct first-order fully decoupled scheme, and prove that the numerical scheme satisfies energy stabilization. For $r\geq 2$, we assume that the solutions to the CHNS model (1.1) in this manuscript exist and satisfy the following regularity assumption:
$\phi\in L^{\infty}\left(0,T;H^{r+1}(\Omega)\right),~{}\phi_{t}\in L^{\infty}\left(0,T;H^{1}(\Omega)\right)\cap L^{2}\left(0,T;H^{1}(\Omega)\right),\phi_{tt}\in L^{2}\left(0,T;L^{2}(\Omega)\right),$ (3.1)
$\mu\in L^{\infty}\left(0,T;H^{r+1}(\Omega)\right),~{}\mu_{t}\in L^{\infty}\left(0,T;H^{1}(\Omega)\right),~{}p\in L^{\infty}\left(0,T;H^{r}(\Omega)\right),~{}p_{t}\in L^{2}\left(0,T;H^{r}(\Omega)\right),$
$\mathbf{u}\in L^{\infty}\left(0,T;\mathbf{H}^{r+1}(\Omega)\right),~{}\mathbf{u}_{t}\in L^{\infty}\left(0,T;\mathbf{H}^{1}(\Omega)\right)\cap L^{2}\left(0,T;\mathbf{H}^{r+1}(\Omega)\right),~{}\mathbf{u}_{tt}\in L^{2}\left(0,T;\mathbf{H}^{-1}(\Omega)\right).$
3.1 Fully discrete scheme
In the subsection, we construct a decoupled, first-order and fully discrete finite elements method based on Euler implicit-explict scheme for solving the CHNS system. For any test functions $w\in H^{1}(\Omega)$, $\varphi\in H^{1}(\Omega)$, $\mathbf{v}\in\mathbf{H}_{0}^{1}(\Omega)$, $q\in L_{0}^{2}(\Omega)$, the weak solutions $\left(\phi,\mu,\mathbf{u},p\right)$ of (1.1) satisfy the following variational forms:
$\left(\partial_{t}\phi,w\right)+M\left(\nabla\mu,\nabla w\right)+b\left(\phi,\mathbf{u},w\right)$ $=0,$ (3.2)
$\left(\mu,\varphi\right)-\lambda\left(\nabla\phi,\nabla\varphi\right)-\frac{\lambda}{\epsilon^{2}}\left(\phi^{3}-\phi,\varphi\right)$ $=0,$ (3.3)
$\left(\partial_{t}\mathbf{u},\mathbf{v}\right)+\nu\left(\nabla\mathbf{u},\nabla\mathbf{v}\right)+\bm{B}\left(\mathbf{u},\mathbf{u},\mathbf{v}\right)+\left(\nabla p,\mathbf{v}\right)-b\left(\phi,\mathbf{v},\mu\right)$ $=0,$ (3.4)
$\left(\nabla\cdot\mathbf{u},q\right)$ $=0.$ (3.5)
Let $N$ be a positive integer and $\left\{t^{n}=n\tau\right\}_{n=0}^{N}$ denotes a uniform partition of the time interval $[0,T]$ with a step size $\tau=T/N$. For $n\geq 0$, choosing the initial conditions $\left(\phi_{h}^{0},\mu_{h}^{0},\mathbf{u}_{h}^{0},p_{h}^{0}\right)$ and $\left(\phi_{h}^{n},\mu_{h}^{n},\mathbf{u}_{h}^{n},p_{h}^{n}\right)$, and updating $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)$, we construct the following fully decoupled, energy stable scheme for the CHNS system.
$\mathbf{Step1.}$ Find $\left(\phi_{h}^{n+1},\mu_{h}^{n+1}\right)\in\left(S_{h}^{r},S_{h}^{r}\right)$ such that
$\left(\delta_{\tau}\phi_{h}^{n+1},w_{h}\right)+b\left(\phi_{h}^{n+1},\mathbf{u}_{h}^{n},w_{h}\right)+M\left(\nabla\mu_{h}^{n+1},\nabla w_{h}\right)=0,$ (3.6)
$\left(\mu_{h}^{n+1},\varphi_{h}\right)-\lambda\left(\nabla\phi_{h}^{n+1},\nabla\varphi_{h}\right)-\frac{\lambda}{\epsilon^{2}}\left(\left(\phi_{h}^{n+1}\right)^{3}-\phi_{h}^{n},\varphi_{h}\right)=0,$ (3.7)
$\partial_{n}\phi_{h}^{n+1}|_{\partial\Omega}=0,~{}\partial_{n}\mu_{h}^{n+1}|_{\partial\Omega}=0.$ (3.8)
$\mathbf{Step2.}$ Find $\tilde{\mathbf{u}}_{h}^{n+1}\in\mathbf{X}_{h}^{r}$ such that
$\left(\frac{\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}}{\tau},\mathbf{v}_{h}\right)+\bm{B}\left(\mathbf{u}_{h}^{n},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{v}_{h}\right)+\nu\left(\nabla\tilde{\mathbf{u}}_{h}^{n+1},\nabla\mathbf{v}_{h}\right)+\left(\nabla p_{h}^{n},\mathbf{v}_{h}\right)-b\left(\phi_{h}^{n+1},\mathbf{v}_{h},\mu_{h}^{n+1}\right)=0,$ (3.9)
$\tilde{\mathbf{u}}_{h}^{n+1}|_{\partial\Omega}=0.$ (3.10)
$\mathbf{Step3.}$ Find $\left(p_{h}^{n+1},\mathbf{u}_{h}^{n+1}\right)\in\left(\mathring{S}_{h}^{r-1},\mathbf{X}_{h}^{r}\right)$ such that
$\frac{\mathbf{u}_{h}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}}{\tau}+\nabla\left(p_{h}^{n+1}-p_{h}^{n}\right)=0,$ (3.11)
$\left(\nabla\cdot\mathbf{u}_{h}^{n+1},q_{h}\right)=0,~{}\mathbf{u}_{h}^{n+1}\cdot\mathbf{n}|_{\partial\Omega}=0,$ (3.12)
hold for all $\left(w_{h},\varphi_{h},\mathbf{v}_{h},q_{h}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$, and $n=0,1,2,\cdots,N-1$, where $\phi_{h}^{0}=R_{h}\phi^{0},~{}\mathbf{u}_{h}^{0}=\mathbf{I}_{h}\mathbf{u}^{0}.$ For simplicity of description, we perform the fully discretization to prove that our numerical scheme satisfies energy stabilization. For the convenience of the following theoretical analysis, we set the parameters $\lambda,M,\nu$ to 1.
3.2 Energy stability
For the above scheme, we can establish the theorem of energy stabilization as follows:
Theorem 3.1.
Let $\phi_{h}^{n}\in S_{h}^{r}$ and $\mathbf{u}_{h}^{n}\in\mathbf{X}_{h}^{r}$ and the initial values $|\phi_{h}^{0}|\leq C_{1}$ and $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})\leq C_{0}$, i.e., the initial energy is bounded. For all time partitions $\tau$ and mesh size $h$ such that $\tau\leq\min\{Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\}$, the decoupled scheme (3.6)-(3.12) is uniquely solvable and satisfies discrete energy law as follows:
$\tilde{E}_{h}^{n+1}-\tilde{E}_{h}^{n}\leq-M\tau\|\nabla\mu_{h}^{n+1}\|^{2}-2\nu\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2},$ (3.13)
where $\tilde{E}_{h}^{n+1}$ is defined by
$\tilde{E}_{h}^{n+1}=\lambda\|\nabla\phi_{h}^{n+1}\|^{2}+\frac{\lambda}{2\epsilon^{2}}\|(\phi_{h}^{n+1})^{2}-1\|^{2}+\|\mathbf{u}_{h}^{n+1}\|^{2}+\tau^{2}\|\nabla p_{h}^{n+1}\|^{2},$ (3.14)
and $C_{0}$ and $C_{1}$ are constant that does depend on the initial value of $\phi$ and $\mathbf{u}$.
Proof.
Letting $w_{h}=2\tau\mu_{h}^{n+1}$ in (3.6), we can obtain
$02\left(\phi_{h}^{n+1}-\phi_{h}^{n},\mu_{h}^{n+1}\right)+2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\mu_{h}^{n+1}\right)+2M\tau\|\nabla\mu_{h}^{n+1}\|^{2}=$ 0 . (3.15)
Letting $\varphi_{h}=-2\left(\phi_{h}^{n+1}-\phi_{h}^{n}\right)$ in (3.7), we can obtain
$-2\left(\mu_{h}^{n+1},\phi_{h}^{n+1}-\phi_{h}^{n}\right)+\lambda\left(\|\nabla\phi_{h}^{n+1}\|^{2}-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.16)
$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$ $=0.$
Letting $\mathbf{v}_{h}=2\tau\tilde{\mathbf{u}}_{h}^{n+1}$ in (3.9), we derive that
$\|\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}-\|\mathbf{u}_{h}^{n}\|^{2}+\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+2\nu\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+2\tau\left(\nabla p_{h}^{n},\tilde{\mathbf{u}}_{h}^{n+1}\right)-2\tau\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}\right)=0.$ (3.17)
Next, we rewrite (3.11) as
$\mathbf{u}_{h}^{n+1}+\tau\nabla p_{h}^{n+1}=\tilde{\mathbf{u}}_{h}^{n+1}+\tau\nabla p_{h}^{n}.$ (3.18)
Then we derive from (3.18) that
$\|\mathbf{u}_{h}^{n+1}\|^{2}-\|\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}-2\tau\left(\nabla p_{h}^{n},\tilde{\mathbf{u}}_{h}^{n+1}\right)+\tau^{2}\left(\|\nabla p_{h}^{n+1}\|^{2}-\|\nabla p_{h}^{n}\|^{2}\right)=0.$ (3.19)
Adding the (3.19) and (3.17), we can obtain
$\|\mathbf{u}_{h}^{n+1}\|^{2}$ $-\|\mathbf{u}_{h}^{n}\|^{2}+\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+2\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\tau^{2}\left(\|\nabla p_{h}^{n+1}\|^{2}-\|\nabla p_{h}^{n}\|^{2}\right)$ (3.20)
$=2\tau\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}\right).$
Combining the (3.15), (3.16) with (3.20), we get
$\lambda\left(\|\nabla\phi_{h}^{n+1}\|^{2}\right.$ $\left.-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.21)
$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$
$+\|\mathbf{u}_{h}^{n+1}\|^{2}-\|\mathbf{u}_{h}^{n}\|^{2}+\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+\tau^{2}\left(\|\nabla p_{h}^{n+1}\|^{2}-\|\nabla p_{h}^{n}\|^{2}\right)$
$=$ $-2M\tau\|\nabla\mu_{h}^{n+1}\|^{2}-2\nu\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+2\tau\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right).$
Recalling inverse inequalities and choosing $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, we have
$2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\bigg{|}\leq$ $~{}2\tau\|\phi_{h}^{n+1}\|_{L^{4}}\|\nabla\mu_{h}^{n+1}\|\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}$ (3.22)
$\leq$ $~{}C\tau\|\phi_{h}^{n+1}-\phi_{h}^{n}+\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2}$
$\leq$ $~{}\frac{1}{2}\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}+\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2},$
where $C$ and $C_{2}$ are constant that does not depend on $\tau$ and $h$. Combining (3.22) with (3.21), and summing up form $n=0$ to $m$, we have
$\tilde{E}_{h}^{m+1}-\tilde{E}_{h}^{0}\leq-\tau\sum_{n=0}^{m}\left(2\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}\right)-\sum_{n=0}^{m}\left(1-\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}.$ (3.23)
In order to the energy to satisfy the law of decay when $m=0$, the following conditions need to be satisfied
$\tau\leq\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}}.$ (3.24)
Since $\|\phi_{h}^{0}\|_{L^{4}}\leq C\|\phi_{h}^{0}\|$ is bounded, the $\tau$ of the above equation (3.24) can be taken to a valid value such that
$\tilde{E}_{h}^{1}-\tilde{E}_{h}^{0}\leq 0.$ (3.25)
So we can obtain that $\|\phi_{h}^{1}\|$ is bounded. When $m=1$, we have
$\tau\leq\min\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}}\bigg{\}},$ (3.26)
then we can dervie that $\|\phi_{h}^{2}\|\leq C$. By analogy, we can obatin the following result
$\tau\leq\min_{0\leq n\leq m}\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},~{}\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}},~{}\cdots,~{}\frac{Ch}{\|\phi_{h}^{n}\|_{L^{4}}^{2}}\bigg{\}}.$ (3.27)
Thus, we can always obtain a positive number $C$ such that $\|\phi_{h}^{n}\|_{L4}\leq C$ is bounded, and derive the desired result (3.13). ∎
Remark 3.1.
The existence of the solution for scheme (3.6)-(3.12) can be proved using the standard argument of the Leray-Schauder fixed point theorem (see [15, 4] for details of similar problems).
Then, we can derive the following lemma regarding the boundedness of the numerical solutions obtained from the scheme (3.6)-(3.12).
Lemma 3.1.
Let $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$ be the unique solution of (3.6)-(3.12). Suppose that $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})<C_{0}$, for any $h,\tau>0$ and $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, the following boundedness hold that
$\max_{0\leq n\leq N-1}\left(\|\mathbf{u}_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}+\frac{\tau}{2\epsilon^{2}}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\tau^{2}\|\nabla p_{h}^{n+1}\|^{2}\right)$ $\leq C_{0},$ (3.28)
$\tau\sum_{n=0}^{N-1}\left(\|\nabla\mu_{h}^{n+1}\|^{2}+\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}\right)$ $\leq C_{0},$ (3.29)
$\sum_{n=0}^{N-1}\left(\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}+\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}\right)$ $\leq C_{0},$ (3.30)
$\frac{\tau^{2}}{\epsilon^{2}}\sum_{n=0}^{N-1}\left[\frac{1}{2}\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right]$ $\leq C_{0},$ (3.31)
where $C_{0}$ is a positive constant dependent on the initial value of $\phi^{0}$ and $\mathbf{u}^{0}$.
The proof of this lemma can be derived from the result of Theorem 3.1, the detailed steps of which we omit.
Lemma 3.2.
For all $m\geq 0$, there exists positive constant $C$ such that
$\|\phi_{h}^{m+1}\|_{H^{1}}^{2}+\tau\sum_{n=0}^{m}\|\mu_{h}^{n+1}\|^{2}\leq C\tau,$ (3.32)
and
$\|\Delta_{h}\phi_{h}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\|\Delta_{h}\mu_{h}^{n+1}\|^{2}\leq C\tau.$ (3.33)
Proof.
Letting $w_{h}=2\tau\phi_{h}^{n+1}$ in (3.6), we obtain
$\|\phi_{h}^{n+1}\|^{2}-\|\phi_{h}^{n}\|^{2}+\|\phi_{h}^{n+1}-\phi_{h}^{n}\|^{2}=-2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\phi_{h}^{n+1}\right)-2\tau\left(\nabla\mu_{h}^{n+1},\nabla\phi_{h}^{n+1}\right).$ (3.34)
Using the Young inequality and Cauchy-Sachwarz inequality, we can obtain
$\big{|}-2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\phi_{h}^{n+1}\right)\big{|}\leq$ $~{}2\tau\|\mathbf{u}_{h}^{n}\|_{L^{4}}\|\nabla\phi_{h}^{n+1}\|\|\phi_{h}^{n+1}\|_{L^{4}}$ (3.35)
$\leq$ $~{}\tau\|\nabla\phi_{h}^{n+1}\|^{2}+C\tau\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\phi_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}\right)$
and
$\big{|}-2\tau\left(\nabla\mu_{h}^{n+1},\nabla\phi_{h}^{n+1}\right)\big{|}\leq\tau\left(\|\nabla\mu_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}\right).$ (3.36)
Combining the above two inequalities with (3.34) and summing up for $n$ from 00 to $m$, we obtain
$\|\phi_{h}^{m+1}\|^{2}\leq\|\phi_{h}^{0}\|^{2}+\tau\sum_{n=0}^{m}\left(\|\nabla\mu_{h}^{n+1}\|^{2}+2\|\nabla\phi_{h}^{n+1}\|^{2}+C\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\phi_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}\right)\right).$ (3.37)
Using the lemma 2.1 and 3.1, we have
$\|\phi_{h}^{m+1}\|^{2}\leq\left(\|\phi_{h}^{0}\|^{2}+c_{1}\left(1+4\tau+c_{2}\right)\right)\exp(c)\leq C\tau.$ (3.38)
Thererfore, we can derive
$\|\phi_{h}^{n+1}\|_{H^{1}}^{2}\leq\|\phi_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}\leq C\tau.$ (3.39)
Testing $\varphi_{h}=\tau\mu_{h}^{n+1}$ in (3.7), we obtain
$\tau\|\mu_{h}^{n+1}\|^{2}=$ $~{}\tau\left(\nabla\phi_{h}^{n+1},\nabla\mu_{h}^{n+1}\right)+\tau\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\mu_{h}^{n+1}\right)$ (3.40)
$\leq$ $~{}\tau\left(\|\nabla\phi_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}+\|\mu_{h}^{n+1}\|^{2}+\|(\phi_{h}^{n+1})^{3}-\phi_{h}^{n}\|^{2}\right).$
Now using the (2.37) in [29], we have
$\|(\phi_{h}^{n+1})^{3}-\phi_{h}^{n}\|^{2}\leq 2\left(\|\phi_{h}^{n+1}\|_{L^{6}}^{6}+\|\phi_{h}^{n}\|^{2}\right)\leq C\left(\|\phi_{h}^{n+1}\|_{H^{1}}^{2}+\|\phi_{h}^{n}\|^{2}\right)\leq C.$ (3.41)
Thus, combining the above inequalities with (3.40) and summing up for $n$ from 00 to $m$, we obtain
$\tau\sum_{n=0}^{m}\|\mu_{h}^{n+1}\|^{2}\leq C\tau.$ (3.42)
From (3.39) and (3.42), we obtain the result of (3.32). We nextly prove the second term of this lemma. Letting $w_{h}=2\tau\Delta_{h}\mu_{h}^{n+1}$ in (3.6), and $\varphi_{h}=\Delta_{h}\delta_{\tau}\phi_{h}^{n+1}$ in (3.7), and combining the results of two terms, we obtain
$~{}\|\Delta_{h}\phi_{h}^{n+1}\|^{2}-\|\Delta_{h}\phi_{h}^{n}\|^{2}+\|\Delta_{h}\left(\phi_{h}^{n+1}-\phi_{h}^{n}\right)\|^{2}+2\tau\|\Delta_{h}\mu_{h}^{n+1}\|^{2}$ (3.43)
$=$ $~{}2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Delta_{h}\mu_{h}^{n+1}\right)+2\tau\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\Delta_{h}\delta_{\tau}\phi_{h}^{n+1}\right).$
Using the Lemma 2.14 in [29], we can estimate the right-hand side of (3.43) as follows:
$\big{|}2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Delta_{h}\mu_{h}^{n+1}\right)\big{|}\leq$ $~{}2\tau\|\mathbf{u}_{h}^{n}\|_{L^{4}}^{2}\|\nabla\phi_{h}^{n+1}\|_{L^{4}}\|\Delta_{h}\mu_{h}^{n+1}\|$ (3.44)
$\leq$ $~{}C\tau\|\nabla\mathbf{u}_{h}^{n}\|\left(\|\nabla\phi_{h}^{n+1}\|+\|\Delta_{h}\phi_{h}^{n+1}\|\right)\|\Delta_{h}\mu_{h}^{n+1}\|$
$\leq$ $~{}C\tau\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\phi_{h}^{n+1}\|^{2}+\|\Delta_{h}\phi_{h}^{n+1}\|^{2}\right)+\tau\|\Delta_{h}\mu_{h}^{n+1}\|^{2}.$
and
$\big{|}2\tau\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\Delta_{h}\delta_{\tau}\phi_{h}^{n+1}\right)\big{|}\leq$ $~{}2\tau\|(\phi_{h}^{n+1})^{3}-\phi_{h}^{n}\|\|\Delta_{h}\delta_{\tau}\phi_{h}^{n+1}\|$ (3.45)
$\leq$ $~{}c\tau^{2}\|\Delta_{h}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+C.$
Combining the above two inequalities (3.44) and (3.45) with (3.43) and summing up for $n$ from 00 to $m$, we obtain
$\|\Delta_{h}\phi_{h}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\|\Delta_{h}\mu_{h}^{n+1}\|^{2}\leq\|\Delta_{h}\phi_{h}^{0}\|^{2}+\tau\sum_{n=0}^{m}\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\phi_{h}^{n+1}\|^{2}+\|\Delta_{h}\phi_{h}^{n+1}\|^{2}\right).$ (3.46)
Using the lemma 2.1 and 3.1, we obtain the result of this lemma. ∎
Remark 3.2.
We note that the bound $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$ in Lemma 3.1 and 3.2 is attainable, since from (4.26) and Lemma 3.2 we obtain $\|\phi_{h}^{n}\|_{L^{4}}\leq C\|\phi_{h}^{n}\|_{H^{1}}\leq C$. More similar details about energy stabilization under this discrete scheme can be found in reference [4].
4 Error analysis
In this section, we establish the error analysis for the scheme (3.6)-(3.12). Firstly, we denote $R_{h}:H^{1}(\Omega)\rightarrow S_{h}^{r}$ as the classic Ritz projection [30],
$\left(\nabla\left(\psi-R_{h}\psi\right),\nabla\varphi_{h}\right)=0,\quad\forall~{}\psi\in H^{1}(\Omega),~{}\varphi_{h}\in S_{h}^{r}.$ (4.1)
Now, for $s\in[2,\infty]$ and $n=1,2,\cdots,N$, from the finite element approximation theory [31], it holds that
$\|\psi-R_{h}\psi\|_{L^{s}}+h\|\psi-R_{h}\psi\|_{W^{1,s}}\leq Ch^{r+1}\|\psi\|_{W^{r+1,s}},$ (4.2)
$\|\psi-R_{h}\psi\|_{H^{-1}}\leq Ch^{r+1}\|\psi\|_{H^{r+1}},$ (4.3)
$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|+h\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (4.4)
$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{-1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (4.5)
where $\delta_{\tau}$ is defined in (2.7). Let $I_{h}:L^{2}(\Omega)\rightarrow S_{h}^{r}$ and $\mathbf{I}_{h}:\mathbf{L}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}$ are defined the $L^{2}$ projection operators [32, 33] by
$\left(v-I_{h}v,w_{h}\right)=0,~{}\forall w_{h}\in S_{h}^{r},$ (4.6)
$\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v},\mathbf{w}_{h}\right)=0,~{}\forall\mathbf{w}_{h}\in\mathbf{X}_{h}^{r}.$ (4.7)
It is known to all that the $L^{2}$ projection satisfies the following estimates,
$\left\|v-I_{h}v\right\|+h\left\|\nabla\left(v-I_{h}v\right)\right\|\leq Ch^{r+1}\|v\|_{H^{r+1}},$ (4.8)
$\left\|\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right\|+h\left\|\nabla\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right)\right\|\leq Ch^{r+1}\|\mathbf{v}\|_{H^{r+1}}.$ (4.9)
For $\left(\mathbf{v}_{h},q_{h}\right)\in\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$, the Stokes projection [34] $P_{h}$ or $\mathbf{P}_{h}:\mathbf{H}_{0}^{1}(\Omega)\times L_{0}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$ are defined by
$\left(\nabla\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),\nabla\mathbf{v}_{h}\right)-\left(p-P_{h}(\mathbf{u},p),\nabla\cdot\mathbf{v}_{h}\right)=0,$ (4.10)
$\left(\nabla\cdot\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),q_{h}\right)=0.$ (4.11)
To be the simplicity of notations, we denote $P_{h}p=P_{h}(\mathbf{u},p)$ and $\mathbf{P}_{h}\mathbf{u}=\mathbf{P}_{h}(\mathbf{u},p)$.
Lemma 4.1 ([35, 25]).
Under the assumption of regularity (3.1), it holds that
$\|\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\|+h\left(\|\nabla\left(\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\right)\|+\|p-P_{h}p\|\right)\leq Ch^{r+1}\left(\|\mathbf{u}\|_{H^{r+1}}+\|p\|_{H^{r}}\right),$ (4.12)
and
$\|\delta_{\tau}\left(\mathbf{u}^{n}-\mathbf{P}_{h}\mathbf{u}^{n}\right)\|\leq$ $~{}Ch^{r+1}\left(\|\delta_{\tau}\mathbf{u}^{n}\|_{H^{r+1}}+\|\delta_{\tau}p^{n}\|\right.$ (4.13)
$~{}\left.+\|\delta_{\tau}\phi^{n}\|_{H^{r+1}}\|\mu^{n}\|_{H^{2}}+\|\phi^{n-1}\|_{H^{r+1}}\|\delta_{\tau}\mu^{n}\|_{H^{2}}\right).$
where $C$ is the constant the only depends on $\Omega$.
The proof of this lemma is similar to Lemma 3.2 of [25]. Here we omit the detailed procedure and refer the interested reader to [25].
Lemma 4.2 ([36]).
The Stokes projection $\mathbf{P}_{h}(\mathbf{u},p)$ is $\mathbf{H}^{1}$ stable in the case such that
$\|\mathbf{P}_{h}(\mathbf{u},p)\|_{H^{1}}\leq C\left(\|\mathbf{u}\|_{H^{2}}+\|p\|_{H^{1}}\right).$ (4.14)
Next we introduce the discrete Laplacian operator [37] $\Delta_{h}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ such that
$\left(-\Delta_{h}v_{h},\xi_{h}\right)=\left(\nabla v_{h},\nabla\xi_{h}\right),$ (4.15)
$\left(\nabla(-\Delta_{h}^{-1})v_{h},\nabla\xi_{h}\right)=\left(v_{h},\xi_{h}\right).$ (4.16)
By defining the discrete Stokes operator $A_{h}=-\mathbf{I}_{h}\Delta_{h}$ [26], it holds
$\|v_{h}\|_{H^{2}}=\|A_{h}v_{h}\|,~{}\|v_{h}\|_{H^{-1}}=\|A_{h}^{-1/2}v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r},$ (4.17)
$\|A_{h}^{1/2}v_{h}\|=\|\nabla v_{h}\|,~{}\|\nabla A_{h}^{-1/2}v_{h}\|=\|v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r}.$ (4.18)
If $v_{h}$ is a constant, we denote $(-\Delta_{h})^{\frac{1}{2}}v_{h}=0$ and $(-\Delta_{h})^{-\frac{1}{2}}v_{h}=0$.
Lemma 4.3 ([25]).
For $v_{h}\in\mathring{S}_{h}^{r-1}$, and the operators $\left(-\Delta_{h}\right)^{1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ and $\left(-\Delta_{h}\right)^{-1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$, we have the estimates
$\|v_{h}\|_{H^{1}}\leq C\|\left(-\Delta_{h}\right)^{1/2}v_{h}\|\leq c\|v_{h}\|_{H^{1}},$ (4.19)
$\|v_{h}\|_{H^{-1}}\leq C\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|\leq c\|v_{h}\|_{H^{-1}},$ (4.20)
and
$\|v_{h}\|^{2}\leq\varepsilon\|\nabla v_{h}\|^{2}+C\varepsilon^{-1}\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|^{2},$ (4.21)
where $\varepsilon$ is an arbitrary positive constant independent of $h$.
Finally, we note the well-known inverse inequalities [25]
$\|v_{h}\|_{L^{4}}\leq Ch^{-1/2}\|v_{h}\|,\quad\forall v_{h}\in S_{h}^{r},$ (4.22)
$\|v_{h}\|_{L^{\infty}}\leq C\sqrt{\ln(1/h)}\|v_{h}\|_{H^{1}},\quad\forall v_{h}\in S_{h}^{r}.$ (4.23)
Lemma 4.4 ([38, 29]).
Suppose $g\in H^{1}(\Omega)$, and $v\in\mathring{S}_{h}^{r-1}$,
$|(g,v)|\leq C\|\nabla g\|\|v\|_{H^{-1}},$ (4.24)
and for all $g\in L_{0}^{2}(\Omega)$, we have
$\|g\|_{H^{-1}}\leq C\|g\|.$ (4.25)
where $C$ is positive constant independent of $\tau$ and $h$.
The proof of the lemma 4.4 is similar to Lemma 3.2 of [38], here we omit it. The following inequalities hold [39]
$\|\phi\|_{L^{r}}\leq C\|\phi\|_{H^{1}},~{}\forall\phi\in H^{1}(\Omega),~{}r\in[2,\infty],$ (4.26)
$\|\phi\|_{L^{\infty}}\leq C\|\phi\|^{\frac{1}{2}}\left(\|\phi\|^{2}+\|\Delta\phi\|^{2}\right)^{\frac{1}{2}},\forall\phi\in H^{2}(\Omega),$ (4.27)
$\|\bm{v}\|_{L^{r}}\leq C\|\nabla\bm{v}\|^{2},~{}(2\leq r\leq 6),\quad\|\bm{v}\|_{L^{4}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|\nabla\bm{v}\|^{\frac{1}{2}},~{}\forall\bm{v}\in\mathbf{H}_{0}^{1}(\Omega),$ (4.28)
$\|\bm{v}\|_{H^{2}}\leq C\|A\bm{v}\|^{2},\quad\|\bm{v}\|_{L^{\infty}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|A\bm{v}\|^{\frac{1}{2}},\forall\bm{v}\in\mathbf{H}^{2}(\Omega)\cap\mathbf{H}_{0}^{1}(\Omega),$ (4.29)
where $A$ is the continuous Stokes operator [34].
And then, we denote
$e_{\phi}^{n+1}=\phi^{n+1}-\phi_{h}^{n+1}=\phi^{n+1}-R_{h}\phi^{n+1}+R_{h}\phi^{n+1}-\phi_{h}^{n+1}=\Phi_{\phi h}^{n+1}+\Theta_{\phi h}^{n+1},$
$e_{\mu}^{n+1}=\mu^{n+1}-\mu_{h}^{n+1}=\mu^{n+1}-R_{h}\mu^{n+1}+R_{h}\mu^{n+1}-\mu_{h}^{n+1}=\Phi_{\mu h}^{n+1}+\Theta_{\mu h}^{n+1},$
and
$e_{\tilde{\mathbf{u}}}^{n+1}=\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\tilde{\mathbf{u}}h}^{n+1},$
$e_{\mathbf{u}}^{n+1}=\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\mathbf{u}h}^{n+1},$
$e_{p}^{n+1}=p^{n+1}-p_{h}^{n+1}=p^{n+1}-P_{h}p^{n+1}+P_{h}p^{n+1}-p_{h}^{n+1}=\Phi_{ph}^{n+1}+\Theta_{ph}^{n+1}.$
For $\left(e_{\phi}^{n+1},e_{\mu}^{n+1},e_{\tilde{\mathbf{u}}}^{n+1},e_{\mathbf{u}}^{n+1},e_{p}^{n+1}\right)$ and the definition of Ritz projection (4.1) and Stokes projection (4.10)-(4.11), we subtract (3.2)-(3.5) from (3.6)-(3.12) to obtain the following error equations,
$\left(\delta_{\tau}e_{\phi}^{n+1},w_{h}\right)+M\left(\nabla e_{\mu}^{n+1},\nabla w_{h}\right)+b\left(\phi^{n+1},\mathbf{u}^{n+1},w_{h}\right)-b\left(\phi_{h}^{n+1},\mathbf{u}_{h}^{n},w_{h}\right)-\left(R_{1}^{n+1},w_{h}\right)$ $=0,$ (4.30)
$\left(e_{\mu}^{n+1},\varphi_{h}\right)-\lambda\left(\nabla e_{\phi}^{n+1},\nabla\varphi_{h}\right)-\frac{\lambda}{\epsilon^{2}}\left((\phi_{h}^{n+1})^{3}-(\phi^{n+1})^{3}-e_{\phi}^{n},\varphi_{h}\right)$ $=0,$ (4.31)
$\left(\frac{e_{\tilde{\mathbf{u}}}^{n+1}-e_{\mathbf{u}}^{n}}{\tau},\mathbf{v}_{h}\right)+\nu\left(\nabla e_{\tilde{\mathbf{u}}}^{n+1},\nabla\mathbf{v}_{h}\right)+\left(\nabla e_{p}^{n},\mathbf{v}_{h}\right)+\left(\bm{B}\left(\mathbf{u}^{n+1},\mathbf{u}^{n+1},\mathbf{v}_{h}\right)-\bm{B}\left(\mathbf{u}_{h}^{n},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{v}_{h}\right)\right)$
$+\left(b\left(\phi_{h}^{n+1},\mathbf{v}_{h},\mu_{h}^{n+1}\right)-b\left(\phi^{n+1},\mathbf{v}_{h},\mu^{n+1}\right)\right)-\left(R_{2}^{n+1},\mathbf{v}_{h}\right)$ $=0,$ (4.32)
$\frac{e_{\mathbf{u}}^{n+1}-e_{\tilde{\mathbf{u}}}^{n+1}}{\tau}+\nabla\left(e_{p}^{n+1}-e_{p}^{n}\right)-R_{3}^{n+1}$ $=0,$ (4.33)
$\left(\nabla\cdot e_{\mathbf{u}}^{n+1},q_{h}\right)$ $=0,$ (4.34)
where the truncation errors $R_{1}^{n+1}$, $R_{2}^{n+1}$ and $R_{3}^{n+1}$ is denoted by
$R_{1}^{n+1}=\delta_{\tau}R_{h}\phi^{n+1}-\partial_{t}\phi^{n+1},R_{2}^{n+1}=\delta_{\tau}\mathbf{P}_{h}\mathbf{u}^{n+1}-\partial_{t}\mathbf{u}^{n+1},R_{3}^{n+1}=\nabla\left(P_{h}p^{n+1}-P_{h}p^{n}\right),$ (4.35)
and the $\delta_{\tau}$ in time scheme is defined by (2.7). Given the above error equations, we aim to derive the optimal $L^{2}$ error estimate for the CHNS system, which is outlined in the following main result.
Theorem 4.1.
Suppose that the model (1.1)-(1.2) has a unique solution $\left(\phi,\mu,\mathbf{u},p\right)$ satisfying the regularities (3.1). Then $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$ is unique solution in the fully discrete scheme (3.6)-(3.12). for a positive constant $\tau_{0}$ such that $0<\tau\leq\tau_{0}$, the numerical solutions satisfy the error estimates as follows:
$\max_{0\leq n\leq N-1}\|\phi^{n+1}-\phi_{h}^{n+1}\|+\left(\tau\sum_{n=0}^{N-1}\|\mu^{n+1}-\mu_{h}^{n+1}\|^{2}\right)^{\frac{1}{2}}\leq C(\tau+h^{r+1}),$ (4.36)
$\max_{0\leq n\leq N-1}\|\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}\|\leq C(\tau+h^{r+1}),$ (4.37)
$\tau\sum_{n=0}^{N-1}\|\nabla\left(\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}\right)\|^{2}\leq C(\tau^{2}+h^{2(r+1)}),$ (4.38)
$\tau\sum_{n=0}^{N-1}\|p^{n+1}-p_{h}^{n+1}\|^{2}\leq C\left(\tau^{2}+h^{2r}\right),$ (4.39)
where $C$ is a positive constant independent of $\tau$ and $h$.
To prove the theorem, we need the following lemmas.
Lemma 4.5.
Under the assumption of regularity (3.1), for any $\tau\geq 0$ and $n=0,1,2,\cdots,N-1$, the truncation errors satisfy
$\|R_{1}^{n+1}\|+\|R_{2}^{n+1}\|+\|R_{3}^{n+1}\|+\|R_{4}^{n+1}\|\leq C\tau.$ (4.40)
The detailed proof of this Lemma can refer to [38, 29].
Lemma 4.6.
For any $h,\tau>0$, we have
$\|\nabla\left((\phi^{n+1})^{3}-(\phi_{h}^{n+1})^{3}\right)\|\leq C\|\nabla e_{\phi}^{n+1}\|.$ (4.41)
where $C$ is independent of $\tau$ and $h$, ${}n=0,1,2,\cdots,N-1.$
Proof.
According to [4, 29], we have
$\|\nabla\left((\phi^{n+1})^{3}-(\phi_{h}^{n+1})^{3}\right)\|$ $\leq 3\|(\phi_{h}^{n+1})^{2}\nabla e_{\phi}^{n+1}\|+3\|\nabla\phi^{n+1}\left(\phi^{n+1}+\phi_{h}^{n+1}\right)e_{\phi}^{n+1}\|$ (4.42)
$\leq 3\|\phi_{h}^{n+1}\|_{L^{\infty}}^{2}\|\nabla e_{\phi}^{n+1}\|+3\|\nabla\phi^{n+1}\|_{L^{6}}\|\phi^{n+1}+\phi_{h}^{n+1}\|_{L^{6}}\|e_{\phi}^{n+1}\|_{L^{6}}$
$\leq 3\left(\|\phi_{h}^{n+1}\|_{L^{\infty}}^{2}+C\|\nabla\phi^{n+1}\|_{L^{6}}\|\phi^{n+1}+\phi_{h}^{n+1}\|_{H^{1}}\right)\|\nabla e_{\phi}^{n+1}\|$
$\leq C\|\nabla e_{\phi}^{n+1}\|.$
Then, using the Lemma 3.1 and the regularity (3.1), we get the result. ∎
Lemma 4.7.
Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
$\|\Theta_{\phi h}^{m+1}\|_{H^{1}}^{2}$ $+\tau\sum_{n=0}^{m}\|\nabla e_{\mu}^{n+1}\|_{H^{1}}^{2}$ (4.43)
$\leq$ $~{}C\left(\tau^{2}+h^{2r}+\tau\sum_{n=0}^{m}\left(\|\phi^{n+1}\|_{H^{2}}^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|\nabla\phi_{h}^{n+1}\|^{2}\right)\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}\right).$
Proof.
To obtain estimates for $e_{\phi}^{n+1}$ and $e_{\mu}^{n+1}$, testing $w_{h}=2\tau\left(\Theta_{\mu h}^{n+1},\Theta_{\phi h}^{n+1}\right)$ and $\varphi_{h}=2\tau\left(\Theta_{\mu h}^{n+1},\delta_{\tau}\Theta_{\phi h}^{n+1}\right)$ in (4.30) and (4.31), we get
$\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}-\|\Theta_{\phi h}^{n}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}-\Theta_{\phi h}^{n}\|_{H^{1}}^{2}+\tau\left(\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}+\|e_{\mu}^{n+1}\|_{H^{1}}^{2}\right)$ (4.44)
$=$ $~{}\tau\|\Phi_{\mu h}\|_{H^{1}}^{2}+2\tau\left(R_{1}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)$
$+2\tau\left((\phi^{n+1})^{3}-\phi^{n}-\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n}\right),\Theta_{\mu h}^{n+1}-\delta_{\tau}\Theta_{\phi h}^{n+1}\right)$
$-2\tau\left(\delta_{\tau}\Phi_{\phi h}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)+2\tau\left(\Phi_{\mu h}^{n+1},\delta_{\tau}\Theta_{\phi h}^{n+1}\right)$
$=$ $~{}\tau\|\Phi_{\mu h}^{n+1}\|_{H^{1}}^{2}+\sum_{n=1}^{5}T_{i}.$
We now estimate the right-hand side terms of (4.44).
$\big{|}T_{1}\big{|}=\bigg{|}2\tau\left(R_{1}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)\bigg{|}\leq\frac{\tau}{7}\left(\|\Theta_{\mu h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}\|_{H^{1}}^{2}\right)+C\tau\|R_{1}^{n+1}\|^{2},$ (4.45)
and
$\big{|}T_{2}\big{|}=$ $~{}\bigg{|}-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)\bigg{|}$ (4.46)
$=$ $~{}2\tau\bigg{|}\left(\left(\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}\right)\cdot\nabla\phi^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)+\left(\mathbf{u}_{h}^{n}\cdot\nabla\left(\phi^{n+1}-\phi_{h}^{n+1}\right),\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)$
$+\left(\left(\mathbf{u}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\cdot\nabla\phi^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)\bigg{|}$
$=$ $~{}2\tau\bigg{|}\left(e_{\mathbf{u}}^{n+1}\cdot\nabla\phi^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)+\left(\mathbf{u}_{h}^{n}\cdot\nabla e_{\phi}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)$
$+\left(\left(\mathbf{u}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\cdot\nabla\phi^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}\frac{\tau}{7}\left(\|\Theta_{\mu h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}\right)+C\tau\|\phi^{n+1}\|_{H^{2}}^{2}\left(\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)$
$+C\tau\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\Phi_{\phi h}^{n+1}\|^{2}+\|\nabla\Theta_{\phi h}^{n+1}\|^{2}\right)+C\tau\left(\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;L^{2}(\Omega))}^{2}\|\nabla\phi^{n+1}\|^{2}\right).$
We can recast (4.30) by the following
$\delta_{\tau}\Theta_{\phi h}^{n+1}=\Delta_{h}e_{\mu}^{n+1}-\delta_{\tau}\Phi_{\phi h}^{n+1}-\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1}\right)+R_{1}^{n+1}.$ (4.47)
Using the (4.26), (4.28), Lemma 4.1 and 4.2, we can obtain
$\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-1}}\leq C\|\nabla e_{\mu}^{n+1}\|+\|R_{1}^{n+1}\|$ (4.48)
$+C\tau\left(\|\phi^{n+1}\|_{H^{2}}+\|\nabla\mathbf{u}_{h}^{n}\|+\|\nabla\mathbf{u}_{h}^{n}\|\|\phi_{h}^{n+1}\|_{H^{1}}\right)\left(\|\Phi_{\mathbf{u}h}^{n+1}\|+\|\Theta_{\mathbf{u}h}^{n+1}\|+\|\Phi_{\phi h}^{n+1}\|_{H^{1}}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}\right).$
According to the Lemma 4.6 and the inequality (3.41), we estimate the right-hand side of (4.44) as follows
$\big{|}T_{3}\big{|}=$ $~{}2\tau\bigg{|}\left((\phi^{n+1})^{3}-\phi^{n}-\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n}\right),\Theta_{\mu h}^{n+1}-\delta_{\tau}\Theta_{\phi h}^{n+1}\right)\bigg{|}$ (4.49)
$\leq$ $~{}C\tau\bigg{|}\left((\phi^{n+1})^{3}-(\phi_{h}^{n+1})^{3},\Theta_{\mu h}^{n+1}-\delta_{\tau}\Theta_{\phi h}^{n+1}\right)+\left(\phi^{n}-\phi_{h}^{n},\Theta_{\mu h}^{n+1}-\delta_{\tau}\Theta_{\phi h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}C\tau\left(\|(\phi^{n+1})^{3}-(\phi_{h}^{n+1})^{3}\|\|\Theta_{\mu h}^{n+1}\|+\|\nabla\left((\phi^{n+1})^{3}-(\phi_{h}^{n+1})^{3}\right)\|\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-1}}\right)$
$+C\tau\left(\|\phi^{n}-\phi_{h}^{n}\|\|\Theta_{\mu h}^{n+1}\|+\|\nabla\left(\phi^{n}-\phi_{h}^{n}\right)\|\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-1}}\right)$
$\leq$ $~{}C\tau\left(\|\Theta_{\mu h}^{n+1}\|^{2}+\|\nabla e_{\mu}^{n+1}\|^{2}+\|R_{1}^{n+1}\|^{2}\right)+C\tau\left(\|\phi^{n+1}\|_{H^{2}}^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right.$
$\left.+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|\phi_{h}^{n+1}\|_{H^{1}}^{2}\right)\left(\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}+\|\Phi_{\phi h}^{n+1}\|_{H^{1}}^{2}\right)+Ch^{2(r+1)}.$
Using the (4.3) and (4.48), we can obtain error estimates of $T_{4}$ and $T_{5}$ as follows:
$\big{|}T_{4}\big{|}=$ $~{}\big{|}-2\tau\left(\delta_{\tau}\Phi_{\phi h}^{n+1},\Theta_{\mu h}^{n+1}+\Theta_{\phi h}^{n+1}\right)\big{|}$ (4.50)
$\leq$ $~{}2\tau\left(\|\Theta_{\mu h}^{n+1}\|_{H^{1}}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}\right)\|\delta_{\tau}\Phi_{\phi h}^{n+1}\|_{H^{-1}}$
$\leq$ $~{}2\tau\left(\|\Theta_{\mu h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}\right)+Ch^{2(r+1)},$
and
$\big{|}T_{5}\big{|}=$ $\big{|}2\tau\left(\Phi_{\mu h}^{n+1},\delta_{\tau}\Theta_{\phi h}^{n+1}\right)\big{|}\leq 2\tau\|\Phi_{\mu h}^{n+1}\|_{H^{1}}\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-1}}$ (4.51)
$\leq$ $~{}C\|\nabla e_{\mu}^{n+1}\|^{2}+C\tau\|R_{1}^{n+1}\|^{2}+C\tau\left(\|\phi^{n+1}\|_{H^{2}}^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right.$
$\left.+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|\phi_{h}^{n+1}\|_{H^{1}}^{2}\right)\left(\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}\right)+Ch^{2(r+1)}.$
Combining the inequalities $T_{1},\cdots,T_{5}$ with (4.44), and summing from $n=0$ to $m$, we can obtain
$\|\Theta_{\phi h}^{m+1}\|_{H^{1}}^{2}+$ $\frac{\tau}{2}\sum_{n=0}^{m}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}\leq\|\Theta_{\phi h}^{0}\|_{H^{1}}^{2}+C\tau\sum_{n=0}^{m}\|R_{1}^{n+1}\|^{2}+C\tau h^{2(r+1)}$ (4.52)
$+C\tau\sum_{n=0}^{m}\left(\|\phi^{n+1}\|_{H^{2}}^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|\phi_{h}^{n+1}\|_{H^{1}}^{2}\right)\left(\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right.$
$\left.\qquad+\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}\right).$
Using the (4.2), and Lemma 4.1, 3.1, 4.5, and 2.1, we can obtain the result of this Lemma. ∎
Lemma 4.8.
Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
$\|\Theta_{\mathbf{u}h}^{m+1}\|^{2}$ $+\tau^{2}\|\nabla\Theta_{ph}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}\leq C\left(\tau^{2}+h^{2r}+\frac{\tau}{4}\sum_{n=0}^{m}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}\right)$ (4.53)
$+C\tau\sum_{n=0}^{m}\left(\|\phi^{n+1}\|_{H^{2}}^{2}+\|\mu_{h}^{n+1}\|_{H^{1}}^{2}+\|\mu_{h}^{n+1}\|_{H^{1}}^{2}\|\nabla\phi_{h}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n+1}\|^{2}\right.$
$\left.+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|\mathbf{u}_{h}^{n}\|^{2}\right)\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}.$
Proof.
Recasting the projection step (4.33), we can obtain
$\Theta_{\mathbf{u}h}^{n+1}+\tau\nabla\Theta_{ph}^{n+1}=\Theta_{\tilde{\mathbf{u}}h}^{n+1}+\tau\nabla\Theta_{ph}^{n}+\tau\left(R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right).$ (4.54)
Taking the inner product of (4.54) with itself on both sides, we can derive
$\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\tau^{2}\|\nabla\Theta_{ph}^{n+1}\|^{2}=$ $~{}\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\tau^{2}\|\nabla\Theta_{ph}^{n}\|^{2}+\tau^{2}\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}$ (4.55)
$+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},\nabla\Theta_{ph}^{n}\right)+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)$
$+2\tau^{2}\left(\nabla\Theta_{ph}^{n},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right).$
Testing $\mathbf{v}_{h}=2\tau\Theta_{\tilde{\mathbf{u}}h}^{n+1}$ in (4), and combining the (4.55) and then using the Stokes projection (4.10), we obtain
$\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}-\|\Theta_{\mathbf{u}h}^{n}\|^{2}+\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}-\Theta_{\mathbf{u}h}^{n}\|^{2}+\tau^{2}\left(\|\Theta_{ph}^{n+1}\|^{2}-\|\Theta_{ph}^{n}\|^{2}\right)+2\tau\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}$ (4.56)
$=$ $2\tau\left(\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)+2\tau^{2}\left(\nabla\Theta_{ph}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},\nabla\left(\Phi_{ph}^{n+1}-\Phi_{ph}^{n}\right)\right)$
$+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)+\tau^{2}\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}+2\tau\left(R_{2}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)$
$+2\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right).$
Now, we need to estimate each term on the right-hand sides of (4.56).
$\big{|}2\tau\left(\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\big{|}\leq\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau\|\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1}\|^{2},$ (4.57)
$\bigg{|}2\tau^{2}\left(\nabla\Theta_{ph}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)\bigg{|}$ (4.58)
$\leq$ $~{}C\tau^{3}\|\nabla\Theta_{ph}^{n+1}\|^{2}+C\tau\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}+\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}$
$\leq$ $~{}C\tau^{3}\|\nabla\Theta_{ph}^{n+1}\|^{2}+\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau\left(\|R_{3}^{n+1}\|^{2}+\tau^{2}\|\nabla\delta_{\tau}\Phi_{ph}^{n+1}\|^{2}\right),$
$\bigg{|}2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},\nabla\left(\Phi_{ph}^{n+1}-\Phi_{ph}^{n}\right)\right)\bigg{|}=$ $~{}\bigg{|}2\tau^{2}\left(\delta_{\tau}\Phi_{ph}^{n+1},\nabla\cdot\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}\leq C\tau^{3}\|\delta_{\tau}\Phi_{ph}^{n+1}\|^{2}+\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2},$ (4.59)
$\bigg{|}\tau^{2}\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}\bigg{|}\leq\tau^{2}\|R_{3}^{n+1}\|^{2}+C\tau^{4}\|\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2},$ (4.60)
$\bigg{|}2\tau\left(R_{2}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}\leq C\tau\|R_{2}^{n+1}\|^{2}+\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2},$ (4.61)
$\bigg{|}2\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}=2\tau\bigg{|}\left(e_{\mu}^{n+1}\nabla\phi^{n+1}+\mu_{h}^{n+1}\nabla e_{\phi}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.62)
$\leq$ $~{}2\tau\bigg{|}\left(e_{\mu}\nabla\phi^{n+1},\Theta_{\mathbf{u}h}^{n+1}+\tau\nabla\left(\Theta_{ph}^{n+1}-\Theta_{ph}^{n}\right)-\tau\left(R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)\right)\bigg{|}$
$+2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla e_{\phi}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}2\tau\|e_{\mu}^{n+1}\|_{L^{4}}\|\nabla\phi^{n+1}\|_{L^{4}}\left(\|\Theta_{\mathbf{u}h}^{n+1}\|+\tau\left(\|\nabla\Theta_{ph}^{n+1}\|+\|\nabla\Theta_{ph}^{n}\|\right)+\tau\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|\right)$
$+2\tau\|\mu_{h}^{n+1}\|_{L^{4}}\|\nabla e_{\phi}^{n+1}\|\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|_{L^{4}}$
$\leq$ $~{}\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\frac{\tau}{4}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}$
$+C\tau^{3}\|\phi^{n+1}\|_{H^{2}}^{2}\left(\|\nabla\Theta_{ph}^{n+1}\|^{2}+\|\nabla\Theta_{ph}^{n}\|^{2}+\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}\right)$
$+C\tau\left(\|\mu_{h}^{n+1}\|_{H^{1}}^{2}+\|\phi^{n+1}\|_{H^{2}}^{2}\right)\left(\|\nabla\Theta_{\phi h}^{n+1}\|^{2}+\|\nabla\Phi_{\phi h}^{n+1}\|^{2}+\|\nabla\Theta_{\mathbf{u}h}^{n+1}\|^{2}\right),$
and
$\bigg{|}-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.63)
$=$ $~{}2\tau\bigg{|}\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}^{n}\cdot\nabla\mathbf{u}^{n}+e_{\mathbf{u}}^{n}\cdot\nabla\mathbf{u}^{n}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\tilde{\mathbf{u}}}^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla\left(\mathbf{u}^{n}-\mathbf{u}^{n+1}\right),\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}.$
According to the Taylor expansion and the inequatilies (4.26)-(4.29), we obtain
$\bigg{|}2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}^{n}\cdot\nabla\mathbf{u}^{n},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.64)
$=$ $~{}2\tau\bigg{|}\left(\left(\mathbf{u}^{n+1}-\mathbf{u}^{n}\right)\cdot\nabla\mathbf{u}^{n+1}+\mathbf{u}^{n}\cdot\nabla\left(\mathbf{u}^{n+1}-\mathbf{u}^{n}\right),\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau^{3}\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\left(\|A_{h}\mathbf{u}^{n+1}\|^{2}+\|A_{h}\mathbf{u}^{n}\|^{2}\right),$
and
$\bigg{|}2\tau\left(e_{\mathbf{u}}^{n}\cdot\nabla\mathbf{u}^{n}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\tilde{\mathbf{u}}}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.65)
$\leq$ $~{}2\tau\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|\left(\|e_{\mathbf{u}}^{n}\|_{L^{4}}\|\nabla\mathbf{u}^{n}\|_{L^{4}}+\|\nabla\mathbf{u}_{h}^{n}\|_{L^{4}}\|e_{\tilde{\mathbf{u}}}^{n+1}\|_{L^{4}}\right)$
$\leq$ $~{}\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau\left(\|\nabla\mathbf{u}^{n}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right)\left(\|\Phi_{\mathbf{u}h}^{n}\|^{2}+\|\Theta_{\mathbf{u}h}^{n}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}\right),$
and
$\bigg{|}2\tau\left(\mathbf{u}_{h}^{n}\cdot\nabla\left(\mathbf{u}^{n}-\mathbf{u}^{n+1}\right),\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}\leq 2\tau\|\mathbf{u}_{h}^{n}\|_{L^{4}}\|\nabla\left(\mathbf{u}^{n}-\mathbf{u}^{n+1}\right)\|_{L^{4}}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|$ (4.66)
$\leq$ $~{}\frac{\tau}{8}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau\|\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}^{n}\|^{2}\right).$
Combining the above inequalities (4.57)-(4.66) with (4.56), and summing up for $n$ from 00 to $m$, we obtain
$\|\Theta_{\mathbf{u}h}^{m+1}\|^{2}+\tau^{2}\|\Theta_{ph}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}$ (4.67)
$\leq$ $~{}\|\Theta_{\mathbf{u}h}^{0}\|^{2}+\tau^{2}\|\Theta_{ph}^{0}\|^{2}+C\tau\sum_{n=0}^{m}\left(\|\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|R_{2}^{n+1}\|^{2}+\|R_{3}^{n+1}\|^{2}\right)$
$+\frac{\tau}{}\sum_{n=0}^{m}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}+C\tau^{3}\sum_{n=0}^{m}\left(\|\nabla\Theta_{ph}^{n+1}\|^{2}+\|\nabla\delta_{\tau}\Phi_{ph}^{n+1}\|^{2}+\|\delta_{\tau}\Phi_{ph}^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\left(\|\mu_{h}^{n+1}\|_{H^{1}}^{2}+\|\phi^{n+1}\|_{H^{2}}^{2}\right)\left(\|\nabla\Theta_{\phi h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|^{2}+\|\nabla\Theta_{\mathbf{u}h}^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\left(\|\nabla\mathbf{u}^{n}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right)\left(\|\Phi_{\mathbf{u}h}^{n}\|^{2}+\|\Theta_{\mathbf{u}h}^{n}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\|\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}^{n}\|^{2}\right)$
$+C\tau^{3}\sum_{n=0}^{m}\|\phi^{n+1}\|_{H^{2}}^{2}\left(\|\nabla\Theta_{ph}^{n+1}\|^{2}+\|\nabla\Theta_{ph}^{n}\|^{2}+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\left(\|A_{h}\mathbf{u}^{n+1}\|^{2}+\|A_{h}\mathbf{u}^{n}\|^{2}\right)\right).$
According to the Lemma 4.1 and 4.2, we can derive
[…]
and
[…]
Finally, using the Lemma 4.1, 3.1, 3.2, 4.5, the inequalities (4.2)-(4.5) and the discrete Grönwall’s lemma 2.1, we can ontain the proof of this lemma. ∎
Therefore, we can obtain the following theorem based on the above lemmas.
Theorem 4.2.
Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
$\|e_{\phi}^{m+1}\|_{H^{1}}^{2}+\|e_{\mathbf{u}}^{m+1}\|^{2}+\tau^{2}\|\nabla e_{p}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\left(\|e_{\mu}^{n+1}\|_{H^{1}}^{2}+\|\nabla e_{\tilde{\mathbf{u}}}^{n+1}\|^{2}\right)\leq C\left(\tau^{2}+h^{2r}\right).$ (4.70)
Next, in order to prove the optimal $L^{2}$ error estimates for the numerical solutions of the fully discrete scheme (3.6)-(3.12), we need to prove the following lemma.
Lemma 4.9.
Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
$\|e_{\phi}^{m+1}\|^{2}+\sum_{n=0}^{m}\|e_{\mu}^{n+1}\|^{2}\leq C\left(\tau^{2}+h^{2(r+1)}\right),$ (4.71)
and
$\|e_{\mathbf{u}}^{m+1}\|^{2}\leq C\left(\tau^{2}+h^{2(r+1)}\right).$ (4.72)
Proof.
Testing $w_{h}=2\tau\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}$, $\varphi_{h}=\delta_{\tau}\Delta_{h}^{-1}\Theta_{\phi h}^{n+1}$ and $\mathbf{v}_{h}=2\tau\Theta_{\tilde{\mathbf{u}}h}^{n+1}$ in the equations (4.30)-(4), and using (4.15)-(4.16) and (4.55), we have
$\|\Theta_{\phi h}^{n+1}\|^{2}-\|\Theta_{\phi h}^{n}\|^{2}+\|\Theta_{\phi h}^{n+1}-\Theta_{\phi h}^{n}\|^{2}+\tau\left(\|\Theta_{\mu h}^{n+1}\|^{2}+\|e_{\mu}^{n+1}\|^{2}\right)$ (4.73)
$=$ $~{}\tau\|\Phi_{\mu h}^{n+1}\|^{2}+2\tau\left(\Phi_{\phi h}^{n+1},\delta_{\tau}\Theta_{\phi h}^{n+1}\right)+2\tau\left(R_{1}^{n+1},\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}\right)$
$+2\tau\left((\phi_{h}^{n+1})^{3}-(\phi^{n+1})^{3}-e_{\phi}^{n},\delta_{\tau}\Delta_{h}^{-1}\Theta_{\phi h}^{n+1}\right)$
$=$ $~{}\tau\|\Phi_{\mu h}^{n+1}\|^{2}+\sum_{i=1}^{3}Tb_{i},$
and
$\|\Theta_{{\mathbf{u}}h}^{n+1}\|^{2}-\|\Theta_{\mathbf{u}h}^{n}\|^{2}+\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}-\Theta_{\mathbf{u}h}^{n}\|^{2}+\tau^{2}\left(\|\nabla\Theta_{ph}^{n+1}\|^{2}-\|\nabla\Theta_{ph}^{n}\|^{2}\right)+2\tau\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}$ (4.74)
$=$ $~{}2\tau\left(\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)+2\tau\left(R_{2}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)+2\tau^{2}\left(\nabla\Theta_{ph}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)$
$+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},\nabla\left(\Phi_{ph}^{n+1}-\Phi_{ph}^{n}\right)\right)+2\tau\left(\Theta_{\tilde{\mathbf{u}}h}^{n+1},R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\right)+\tau^{2}\|R_{3}^{n+1}-\tau\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}$
$+2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}\right)$
$+2\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)$
$=$ $~{}\sum_{i=4}^{12}Tb_{i}.$
We now need to estimate the right-hand side terms $Tb_{1},\cdots,Tb_{12}$ of (4.73).
$\bigg{|}Tb_{1}+Tb_{2}\bigg{|}=$ $~{}\bigg{|}2\tau\left(\Phi_{\phi h}^{n+1},\delta_{\tau}\Theta_{\phi h}^{n+1}\right)+2\tau\left(R_{1}^{n+1},\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}\right)\bigg{|}$ (4.75)
$\leq$ $~{}C\tau\left(\|\Phi_{\phi h}^{n+1}\|^{2}+\tau\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|^{2}+\frac{1}{2}\|R_{1}^{n+1}\|^{2}+\frac{1}{2}\|\Theta_{\mu h}^{n+1}\|^{2}\right).$
According to the inequalities (4.47), (4.3) and the fact that $\|\cdot\|_{H^{-2}}\leq\|\cdot\|_{H^{-1}}$, we obtain
$\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-2}}\leq$ $~{}\|\Delta_{h}e_{\mu}^{n+1}\|_{H^{-2}}+\|\delta_{\tau}\Phi_{\phi h}^{n+1}\|_{H^{-1}}+\|\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1}\|_{H^{-1}}$ (4.76)
$\leq$ $~{}\tau\|e_{\mu}^{n+1}\|+\frac{C\tau}{2}\|R_{1}^{n+1}\|+C\tau\left(\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}\|\nabla\phi^{n+1}\|\right)$
$+C\tau\|\nabla\phi^{n+1}\|\left(\|\Phi_{\mathbf{u}h}^{n+1}\|+\|\Theta_{\mathbf{u}h}^{n+1}\|\right)+C\tau\|\nabla\mathbf{u}_{h}^{n+1}\|\left(\|\Phi_{\phi h}^{n+1}\|+\|\Theta_{\phi h}^{n+1}\|\right).$
Using the Lemma 4.6, the inequalities (3.41) and (4.49), the term $Tb_{9}$ is estimated by
$\bigg{|}Tb_{3}\bigg{|}=$ $~{}\bigg{|}2\tau\left((\phi_{h}^{n+1})^{3}-(\phi^{n+1})^{3}-e_{\phi}^{n},\delta_{\tau}\Delta_{h}^{-1}\Theta_{\phi h}^{n+1}\right)\bigg{|}$ (4.77)
$\leq$ $~{}2\tau\bigg{|}\left(\phi_{h}^{n+1})^{3}-(\phi^{n+1})^{3},\delta_{\tau}\Delta_{h}^{-1}\Theta_{\phi h}^{n+1}\right)\bigg{|}+2\tau\bigg{|}\left(e_{\phi}^{n},\delta_{\tau}\Delta_{h}^{-1}\Theta_{\phi h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}C\left(\|\phi_{h}^{n+1})^{3}-(\phi^{n+1})^{3}\|+\|e_{\phi}^{n}\|\right)\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-2}}$
$\leq$ $~{}\frac{\tau}{2}\|e_{\mu}^{n+1}\|^{2}+\frac{C\tau}{2}\|R_{1}^{n+1}\|^{2}+C\tau^{3}\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\|\nabla\phi^{n+1}\|^{2}$
$+C\tau\left(\|\nabla\phi^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}+\|\Theta_{\phi h}^{n}\|^{2}+\|\Phi_{\phi h}^{n}\|^{2}\right)\left(\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right.$
$\left.+\|\Theta_{\phi h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|^{2}\right)+Ch^{2(r+1)}.$
Using a similar proof to that of the Lemma 4.8, it can easily be proved that
$\bigg{|}\sum_{i=4}^{9}Tb_{i}\bigg{|}\leq$ $~{}C\tau\left(\frac{1}{3}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\|\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|R_{2}^{n+1}\|^{2}\right)$ (4.78)
$+C\tau\left(\tau^{2}\|\nabla\Theta_{ph}^{n+1}\|^{2}+\|R_{3}^{n+1}\|^{2}+\tau^{3}\|\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}\right)$
$+C\tau^{3}\left(\|\delta_{\tau}\Phi_{ph}^{n+1}\|_{H^{1}}^{2}+\|\delta_{\tau}\nabla\Phi_{ph}^{n+1}\|^{2}\right).$
We next estimate the three terms $Tb_{10}$, $Tb_{11}$ and $Tb_{12}$ as follows
$\bigg{|}Tb_{10}\bigg{|}=$ $~{}\bigg{|}2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1},\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}\right)\bigg{|}$ (4.79)
$\leq$ $~{}2\tau\left(e_{\mathbf{u}}^{n+1}\cdot\nabla\phi^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\phi}^{n+1}+\left(\mathbf{u}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\cdot\nabla\phi^{n+1},\Delta_{h}^{-1}\Theta_{\mu h}^{n+1}\right)$
$\leq$ $~{}\frac{C\tau}{4}\|\Theta_{\mu h}^{n+1}\|^{2}+C\tau\left(\|\phi^{n+1}\|_{H^{1}}^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\|\phi^{n+1}\|_{H^{1}}^{2}\right),$
and
$\bigg{|}Tb_{11}\bigg{|}=$ $~{}\bigg{|}2\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.80)
$=$ $~{}2\tau\bigg{|}\left(e_{\mu}^{n+1}\nabla\phi^{n+1}+\mu_{h}^{n+1}\nabla e_{\phi}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$
$\leq$ $~{}2\tau\|e_{\mu}^{n+1}\|_{L^{4}}\|\phi^{n+1}\|_{L^{4}}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|+2\tau\|\mu_{h}^{n+1}\|_{L^{4}}\|e_{\phi}^{n+1}\|_{L^{4}}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|$
$\leq$ $~{}\frac{\tau}{3}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\frac{\tau}{4}\|e_{\mu}^{n+1}\|^{2}+C\tau^{3}\left(\|\mu_{h}^{n+1}\|_{H^{1}}^{2}+\|\phi^{n+1}\|_{H^{1}}^{2}\right)\left(\|\Phi_{\phi h}^{n+1}\|^{2}+\|\Theta_{\phi h}^{n+1}\|^{2}\right),$
and
$\bigg{|}Tb_{12}\bigg{|}=$ $~{}\bigg{|}-2\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)\bigg{|}$ (4.81)
$=$ $~{}2\tau\left(e_{\mathbf{u}}^{n+1}\cdot\nabla\mathbf{u}^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\mathbf{u}}^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\tilde{\mathbf{u}}}^{n+1}+\left(\mathbf{u}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\cdot\nabla\mathbf{u}^{n+1},\Theta_{\tilde{\mathbf{u}}h}^{n+1}\right)$
$\leq$ $~{}\frac{\tau}{3}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+C\tau\left(\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right)\left(\|\Theta_{{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)$
$+C\tau\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right).$
Combining the above inequalities $Tb_{1},\cdots,Tb_{12}$ and summing up from $n=0$ to $m$, we have
$\|\Theta_{\phi h}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\left(\frac{1}{2}\|\Theta_{\mu h}^{n+1}\|^{2}+\frac{1}{2}\|e_{\mu}^{n+1}\|^{2}\right)$ (4.82)
$\leq$ $~{}\|\Theta_{\phi h}^{0}\|^{2}+\tau\sum_{n=0}^{m}\|\Phi_{\mu h}^{n+1}\|^{2}+C\tau\sum_{n=0}^{m}\left(\|\Phi_{\phi h}^{n+1}\|^{2}+\|R_{1}^{n+1}\|^{2}+\tau^{2}\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;L^{2}(\Omega))}^{2}\|\nabla\phi^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\left(\|\nabla\phi^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}+\|\Theta_{\phi h}^{n}\|^{2}+\|\Phi_{\phi h}^{n}\|^{2}\right)\left(\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right.$
$\left.+\|\Theta_{\phi h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|^{2}\right)+Ch^{2(r+1)},$
and
$\|\Theta_{{\mathbf{u}}h}^{m+1}\|^{2}+\tau^{2}\|\Theta_{ph}^{m+1}\|+\tau\sum_{n=0}^{m}\|\nabla\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}$ (4.83)
$\leq$ $~{}\|\Theta_{{\mathbf{u}}h}^{0}\|^{2}+\tau^{2}\|\nabla\Theta_{ph}^{0}\|^{2}+C\tau\sum_{n=0}^{m}\left(\|\delta_{\tau}\Phi_{\mathbf{u}h}^{n+1}\|^{2}+\|\Theta_{\mu h}^{n+1}\|^{2}+\|R_{2}^{n+1}\|^{2}+\|R_{3}^{n+1}\|^{2}\right)$
$+C\tau^{3}\sum_{n=0}^{m}\left(\|\mu_{h}^{n+1}\|_{H^{1}}^{2}+\|\phi^{n+1}\|_{H^{1}}^{2}\right)\left(\|\Phi_{\phi h}^{n+1}\|^{2}+\|\Theta_{\phi h}^{n+1}\|^{2}\right)$
$+C\tau^{3}\sum_{n=0}^{m}\left(\|\delta_{\tau}\Phi_{ph}^{n+1}\|^{2}+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\|\nabla\phi^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\left(\left(\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right)\left(\|\Theta_{{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)\right).$
According to the triangle inequality, the inequality (4.3), the Lemma 4.1, 3.1 and 3.2, this completes the proof of Lemma 4.9. ∎
Lemma 4.10.
Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
$\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}\leq C\tau\left(\tau^{2}+h^{2r}\right).$ (4.84)
Proof.
Adding the (4)-(4.33), we obtain
$\left(\delta_{\tau}e_{\mathbf{u}}^{n+1},\mathbf{v}_{h}\right)+\left(\nabla e_{\tilde{\mathbf{u}}}^{n+1},\nabla\mathbf{v}_{h}\right)-\left(e_{p}^{n+1},\nabla\cdot\mathbf{v}_{h}\right)-\left(R_{2}^{n+1},\mathbf{v}_{h}\right)-\left(R_{3}^{n+1},\mathbf{v}_{h}\right)$ (4.85)
$+\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{v}_{h}\right)-\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\phi_{h}^{n+1},\mathbf{v}_{h}\right)$ $=0.$
Testing $\mathbf{v}_{h}=\tau A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})$ and using the (4.17)-(4.18), (4.34), (4.54), we have
$\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}=$ $~{}\tau\left(\Delta_{h}e_{\tilde{\mathbf{u}}}^{n+1},A_{h}^{-1}\left(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)\right)+\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\phi_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)$ (4.86)
$-\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)$
$+\tau\left(R_{2}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)+\tau\left(R_{3}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right).$
According to (4.6) and (3.11), we can obtain $\mathbf{u}_{h}^{n+1}=\mathbf{I}_{h}\tilde{\mathbf{u}}_{h}^{n+1}$. Using the (4.15) and the above results, we have
$\tau\left(\Delta_{h}e_{\tilde{\mathbf{u}}}^{n+1},A_{h}^{-1}\left(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)\right)=$ $~{}-\left(e_{\tilde{\mathbf{u}}}^{n+1},-\tau\Delta_{h}A_{h}^{-1}\left(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)\right)$ (4.87)
$=$ $~{}-\tau\left(e_{{\mathbf{u}}}^{n+1},e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)$
$=$ $~{}-\frac{\tau}{2}\left(\|e_{\mathbf{u}}^{n+1}\|^{2}-\|e_{\mathbf{u}}^{n}\|^{2}+\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|^{2}\right).$
By substituting (4.87) into (4.86), we obtain
$\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}+\frac{\tau}{2}\left(\|e_{\mathbf{u}}^{n+1}\|^{2}-\|e_{\mathbf{u}}^{n}\|^{2}+\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|^{2}\right)$ (4.88)
$=$ $~{}\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\phi_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)$
$-\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)$
$+\tau\left(R_{2}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)+\tau\left(R_{3}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right).$
We then estimate the right-hand side terms of (4.88) as follows
$~{}\bigg{|}\tau\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)\bigg{|}$ (4.89)
$=$ $~{}\bigg{|}\tau\left(e_{\mu}^{n+1}\nabla\phi^{n+1}+\mu_{h}^{n+1}\nabla e_{\phi}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)\bigg{|}$
$\leq$ $~{}\frac{1}{6}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}+C\tau^{2}\left(\|\phi^{n+1}\|_{H^{1}}^{2}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}+\|\mu_{h}^{n+1}\|_{H^{1}}^{2}\|e_{\phi}^{n+1}\|_{H^{1}}^{2}\right),$
and
$\bigg{|}-\tau\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)\bigg{|}$ (4.90)
$=$ $~{}\bigg{|}\tau\left(e_{\mathbf{u}}^{n}\cdot\nabla\mathbf{u}^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\mathbf{u}}^{n+1}+\mathbf{u}_{h}^{n}\cdot\nabla e_{\tilde{\mathbf{u}}}^{n+1}+\left(\mathbf{u}^{n+1}-\mathbf{u}^{n}\right)\cdot\nabla\mathbf{u}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)\bigg{|}$
$\leq$ $~{}\frac{1}{6}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}+C\tau\left(\|e_{\mathbf{u}}^{n}\|^{2}\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|e_{\mathbf{u}}^{n+1}\|^{2}\right)+C\tau^{4}\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;H^{1}(\Omega))}^{2}\|\nabla\mathbf{u}^{n+1}\|^{2},$
and
$\bigg{|}\tau\left(R_{2}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)+\tau\left(R_{3}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)\bigg{|}$ (4.91)
$\leq$ $~{}\frac{1}{6}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}+C\tau^{2}\left(\|R_{2}^{n+1}\|+\|R_{3}^{n+1}\|^{2}\right).$
Combining (4.89)-(4.91) with (4.88) and summing over $n,~{}(0\leq n\leq m)$, we have
$~{}\frac{1}{2}\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}+\frac{\tau}{2}\|e_{\mathbf{u}}^{m+1}\|^{2}$ (4.92)
$\leq$ $~{}\frac{\tau}{2}\|e_{\mathbf{u}}^{0}\|^{2}+C\tau^{2}\sum_{n=0}^{m}\left(\|\phi^{n+1}\|_{H^{1}}^{2}\|e_{\mu}^{n+1}\|_{H^{1}}^{2}+\|\mu_{h}^{n+1}\|_{H^{1}}^{2}\|e_{\phi}^{n+1}\|_{H^{1}}^{2}+\|R_{2}^{n+1}\|+\|R_{3}^{n+1}\|^{2}\right)$
$+C\tau\sum_{n=0}^{m}\left(\|e_{\mathbf{u}}^{n}\|^{2}\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\|e_{\mathbf{u}}^{n+1}\|^{2}\right)+C\tau^{4}\sum_{n=0}^{m}\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\|\nabla\mathbf{u}^{n+1}\|^{2}.$
The proof of this lemma is completed. ∎
Next, we prove the optimal $L^{2}$ error estimate for the pressure. Using the inf-sup condition (2.5) and the equation (4.85), we obtain
$C\|e_{p}^{n+1}\|\leq$ $~{}\frac{\left(e_{p}^{n+1},\nabla\cdot\mathbf{v}_{h}\right)}{\|\nabla\mathbf{v}_{h}\|}$ (4.93)
$\leq$ $~{}[\left(\delta_{\tau}e_{\mathbf{u}}^{n+1},\mathbf{v}_{h}\right)+\left(\nabla e_{\tilde{\mathbf{u}}}^{n+1},\nabla\mathbf{v}_{h}\right)-\left(R_{2}^{n+1},\mathbf{v}_{h}\right)-\left(R_{3}^{n+1},\mathbf{v}_{h}\right)$
$+\left(\mathbf{u}^{n+1}\cdot\nabla\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{v}_{h}\right)-\left(\mu^{n+1}\nabla\phi^{n+1}-\mu_{h}^{n+1}\phi_{h}^{n+1},\mathbf{v}_{h}\right)]/\|\nabla\mathbf{v}_{h}\|$
$\leq$ $~{}\frac{1}{\tau}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}+\|\nabla e_{\tilde{\mathbf{u}}}^{n+1}\|+\|R_{2}^{n+1}\|+\|R_{3}^{n+1}\|$
$+\|\nabla\mathbf{u}^{n+1}\|\left(\|e_{\mathbf{u}}^{n}\|+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}\right)+\|e_{\mu}^{n+1}\|_{H^{1}}\|\nabla\phi^{n+1}\|+\|\mu_{h}^{n+1}\|_{H^{1}}\|\nabla e_{\phi}^{n+1}\|.$
Squaring both ends of the (4.93) and multiplying the result by $\tau$, summing over $n~{}(0\leq n\leq m)$, and using the Lemma 3.1, 3.2, 4.5, 4.10 and the Theorem 4.2, we can obtain the result of this lemma as follows
$\tau\sum_{n=0}^{m}\|e_{p}^{n+1}\|^{2}\leq C\left(\tau^{2}+h^{2r}\right).$ (4.94)
Up to this point, by utilizing the above lemmas and theorems, we have successfully arrived at the result of Theorem 4.1.
5 Numerical experiments
In this section, we conduct three numerical experiments for the following purposes: 1) to validate the convergence rates as stated in Theorem 4.1; 2) to simulate the coarsening dynamics using a random initial phase function; and 3) to demonstrate numerically the unconditional stability of our scheme, which is only conditionally stable in theory. In all tests, we employ the finite element spaces $S^{2}_{h}\times S^{2}_{h}\times\mathbf{X}^{2}_{h}\times S^{1}_{h}$ for the phase field function, chemical potential, velocity, and pressure, respectively; in other words, we use the $P_{2}$ function space for both $\phi^{n+1}_{h}$ and $\mu^{n+1}_{h}$, and the $\mathbf{P}_{2}\times P_{1}$ space for $(\mathbf{u}^{n+1}_{h},p^{n+1}_{h})$.
5.1 Convergence tests
The first test is performed on $\Omega=[0,1]\times[0,1]$ and time $T=0.01$, and the physical parameters are selected as $M=0.1$, $\lambda=0.04$, $\epsilon=0.04$ and $\nu=0.1$. The time-step $\tau$ is set up as $\tau=0.1h^{3}$ to balance the convergence rates between time and space. The manufactured solutions are chosen from [40] and takes the form
$\phi(t,x,y)&=2+\sin(t)\cos(\pi x)\cos(\pi y),\\ \mathbf{u}(t,x,y)&=\left[\pi\sin(\pi x)^{2}\sin(2\pi y)\sin(t),-\pi\sin(\pi y)^{2}\sin(2\pi x)\sin(t)\right]^{\text{T}},\\ p(t,x,y)&=\cos(\pi x)\sin(\pi y)\sin(t),\end{aligned}\right.$ (5.1)
and the exact chemical potential $\mu(t,x,y)$ is obtained by its definition, i.e.,
$\mu(t,x,y):=$ $-\lambda\Delta\phi(t,x,y)+\lambda F^{\prime}\big{(}\phi(t,x,y)\big{)}$ (5.2)
$=$ $-\lambda\Delta\phi(t,x,y)+\frac{\lambda}{\epsilon^{2}}\left(\phi(t,x,y)^{3}-\phi(t,x,y)\right)$
$=$ $~{}2\lambda\pi^{2}\cos(\pi x)\cos(\pi y)\sin(t)$
$-\frac{\lambda}{\epsilon^{2}}\left[\cos(\pi x)\cos(\pi y)\sin(t)-\big{(}\cos(\pi x)\cos(\pi y)\sin(t)+2\big{)}^{3}+2\right].$
For simplicity of notations, we adopt the following notations:
$\|\cdot\|_{\ell^{\infty}(L^{2})}:=\max\limits_{0\leq n\leq N-1}\|\cdot\|_{L^{2}},\quad|\cdot|_{\ell^{\infty}}:=\max\limits_{0\leq n\leq N-1}|\cdot|,\quad\|\cdot\|_{\ell^{2}(L^{2})}:=\sqrt{\tau\sum_{n=0}^{N-1}\|\cdot\|_{L^{2}}}.$ (5.3)
Therefore, according to Theorem 4.1, the following theoretically optimal orders should be observed:
$\left\|\phi-\phi_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\approx\mathcal{O}(h^{3}),\qquad\left\|\mu-\mu_{h}\right\|_{\ell^{2}(L^{2})}\approx\mathcal{O}(h^{3}),$ (5.4)
$\left\|\bm{u}-\bm{u}_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\approx\mathcal{O}(h^{3}),\qquad\left\|p-p_{h}\right\|_{\ell^{2}(L^{2})}\approx\mathcal{O}(h^{2}),$
$\left\|\nabla(\bm{u}-\bm{u}_{h})\right\|_{\ell^{2}(L^{2})}$ $\approx\mathcal{O}(h^{2}).$
Table 1 and Table 2 present the errors and convergence rates, where $L^{2}$ errors and convergence orders, and $H^{1}$ errors and convergence orders for velocity of CHNS are presented by spatial ones after setting $\tau=\mathcal{O}(h^{3})$. From those tables, we can conclude that the numerical results are consistent with the above a prior rates, and hence the scheme proposed here can indeed obtain the optimal convergence rates.
Table 1: $L^{2}$ errors and convergence orders.
$h$ $\left\|\phi-\phi_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\left\|\mu-\mu_{h}\right\|_{\ell^{2}(L^{2})}$ $\left\|p-p_{h}\right\|_{\ell^{2}(L^{2})}$
error rate error rate error rate
1/4 8.5456e--05 - 6.5126e--06 - 8.4150e--03 -
1/8 1.0599e--05 3.0112 8.3148e--07 2.9695 1.7367e--03 3.2502
1/16 1.2872e--06 3.0417 1.0217e--07 3.0247 1.7543e--04 3.1413
1/32 1.5815e--07 3.0249 1.2640e--08 3.0149 2.1352e--05 3.0409
1/64 1.9585e--08 3.0135 1.5707e--08 3.0085 2.6588e--06 3.0074
Table 2: $L^{2}$- and $H^{1}$-norm errors and convergence orders for velocity and time convergence orders for CHNS (presented by spatial orders).
$h$ $\left\|\bm{u}-\bm{u}_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\left\|\nabla(\bm{u}-\bm{u}_{h})\right\|_{\ell^{2}(L^{2})}$
error rate error rate
1/4  3.4875e--04 -  2.1176e--05  -
1/8  7.0458e--05 2.3079  3.9048e--06  2.4395
1/16  8.6735e--06 3.0217  4.9643e--07  2.9754
1/32  1.0212e--06 3.0861  5.8912e--08  3.0747
1/64  1.2489e--07 3.0315  7.2101e--09  3.0305
5.2 Coarsening dynamics
In this example, we employ the Cahn–Hilliard–Navier–Stokes equations (1.1) to simulate the coarsening dynamics, with a random phase function. We set the domain as $\Omega=[0,1]\times[0,1]$, and choose the parameters as $M=0.01$, $\lambda=0.02$, $\epsilon=0.01$, and $\nu=1$, along with a random initial condition for the phase function within the range $[-0.1,0.1]$. The spatial discretization is set at $h=1/64$, and the time step is $\tau=0.001$. We conduct this test up to $T=10$ and record the phase function snapshots at $t=0,0.01,0.05,0.1,0.5,1,3,5,10s$, which are displayed in Figure 1.
Figure 1: Snapshots of phase function at difference times from left to right row by row with $t=0,0.01,0.05,0.1,0.5,1,3,5,10s$, respectively.
5.3 Shape Relaxation with Rotational Boundary Condition
In this section, we present an example of shape relaxation within the domain $\Omega=(0,1)\times(0,1)$, where the boundary condition is set to $\mathbf{u}=(y-0.5,-x+0.5)$ on $\partial\Omega$. We evaluate the performance of the scheme described by (3.6)-(3.12) using critical phase field initial conditions. Specifically, we define $\phi^{0}=1$ within a polygonal subdomain featuring reentrant corners and $\phi^{0}=-1$ in the rest of $\Omega$. The initial velocity is given by $\mathbf{u}^{0}=(y-0.5,-x+0.5)$. This problem has been numerically studied in [4]. For our simulations, we adopt the following parameters: $M=0.001$, $\lambda=0.1$, $\epsilon=0.01$, $\nu=1$, and $T=0.5$. We solve the problem using scheme (3.6)-(3.12) with a $P_{2}$ element for $\phi_{h}^{n+1}$ and $\mu_{h}^{n+1}$, and a $P_{2}\times P_{1}$ element for $\mathbf{u}_{h}^{n+1}$ and $p_{h}^{n+1}$. We run the simulation up to $T=0.5$ and record the phase function snapshots at $t=0,0.01,0.05,0.08,0.1,0.2,0.3,0.5$, which are depicted in Figure 2. Our scheme is unconditionally energy stable in this example, as shown in Figure 3.
Figure 2: Snapshots of phase function at difference times from left to right row by row with $t=0,0.01,0.05,0.08,0.1,0.2,0.3,0.5$, respectively.
5.4 Comparison of time efficiency
In the subsection, we compare it with the following scheme [25] in the computation of the total time,
$\left(\delta_{\tau}\phi_{h}^{n+1},w_{h}\right)+\left(\nabla\mu_{h}^{n+1},\nabla w_{h}\right)+b\left(\phi_{h}^{n},\mathbf{u}_{h}^{n+1},w_{h}\right)={}$ $0,$ (5.5)
$\left(\nabla\phi_{h}^{n+1},\nabla\varphi_{h}\right)-\left(\mu_{h}^{n+1},\varphi_{h}\right)+\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\varphi_{h}\right)={}$ $0,$
$\left(\delta_{\tau}\mathbf{u}_{h}^{n+1},\mathbf{v}_{h}\right)+\left(\nabla\mathbf{u}_{h}^{n+1},\nabla\mathbf{v}_{h}\right)+\bm{B}\left(\mathbf{u}_{h}^{n},\mathbf{u}_{h}^{n+1},\mathbf{v}_{h}\right)-\left(p_{h}^{n+1},\nabla\cdot\mathbf{v}_{h}\right)-b\left(\phi_{h}^{n},\mathbf{v}_{h},\mu_{h}^{n+1}\right)={}$ $0,$
$\left(\nabla\cdot\mathbf{u}_{h}^{n+1},q_{h}\right)={}$ $0$
By using the above scheme and our proposed scheme to the convergence tests of subsection 5.1, we obtain the time results in Table 3. The data in Table 3 shows that our algorithm takes 8.22 hour and the algorithm in [25] takes 23.18 hour, which is about 3 times faster in time than the scheme of [25]. Thus our scheme is more time efficient.
Table 3: Comparison of time consumed in different schemes.
    Scheme     CPU Time (hour)
Cai et al. [25] 23.18
Our scheme 8.22
5.5 Lid driven cavity boundary condition
In the experiment, we take $\Omega=[0,1]^{2}$, $\epsilon=0.02$, $M=0.025$, $\nu=0.1$, $\lambda=0.001$, boundary condition on the top is $g=[1,0]$, initial condition for velocity is $\mathbf{u}_{0}=[1,0]$, $\tau=0.001,h=1/64$. We solve the problem using scheme (3.6)-(3.12) with the $P_{2}\times P_{2}\times P_{2}\times P_{1}$ for $\phi_{h}^{n+1}$, $\mu_{h}^{n+1}$, $\mathbf{u}_{h}^{n+1}$ and $p_{h}^{n+1}$, snapshots time $t=0,5,10,20,30,50$. The evolution of the concentration for this problem is shown through Figure 4 (see [41] for a similar problem).
Figure 3: Energy decay of our scheme in this example.
Figure 4: Evolution of the concentration with different time-step $\tau$.
5.6 Verification of energy stability
In this final test, we aim to verify the energy stability of the new scheme. We use the same settings as those in the coarsening dynamics test (see subsection 5.2), but with varying time steps $\tau$. As shown in Figure 5, our scheme demonstrates unconditional stability with respect to energy dissipation.
Figure 5: Evolution of the energy with different time-step $\tau$.
6 Conclusions
In this work, we propose a decoupling method for the Cahn–Hilliard–Navier–Stokes (CHNS) model, a highly coupled nonlinear system. The method, based on a convex-splitting approach, is first-order, fully discrete, and energy stable, making it efficient and easy to implement. Moreover, we have conducted a rigorous error analysis for the fully discrete scheme and derived optimal error estimates for all relevant functions in the $L^{2}$ norm. The presented numerical experiments are designed to verify the theoretical results of our method. It has been experimentally proven that our scheme is unconditionally stable, a property that has not been theoretically proven and requires further research.
Acknowledgments
This research is supported by the National Natural Science Foundation of China (No.11971337).

### v2

@@
 In an abuse of notation, we use $\bm{g}^{n+1}$ hereafter to denote the value of the exact solution $\bm{g}$ at $t^{n+1}$.
+Next, we denote $R_{h}:H^{1}(\Omega)\rightarrow S_{h}^{r}$ as the classic Ritz projection [30],
+$\left(\nabla\left(\psi-R_{h}\psi\right),\nabla\varphi_{h}\right)=0,\quad\forall~{}\psi\in H^{1}(\Omega),~{}\varphi_{h}\in S_{h}^{r}.$ (2.8)
+Now, for $s\in[2,\infty]$ and $n=1,2,\cdots,N$, from the finite element approximation theory [31], it holds that
+$\|\psi-R_{h}\psi\|_{L^{s}}+h\|\psi-R_{h}\psi\|_{W^{1,s}}\leq Ch^{r+1}\|\psi\|_{W^{r+1,s}},$ (2.9)
+$\|\psi-R_{h}\psi\|_{H^{-1}}\leq Ch^{r+2}\|\psi\|_{H^{r+1}},$ (2.10)
+$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|+h\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (2.11)
+$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{-1}}\leq Ch^{r+2}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (2.12)
+where $\delta_{\tau}$ is defined in (2.7).
+Let $I_{h}:L^{2}(\Omega)\rightarrow S_{h}^{r}$ and $\bm{I}_{h}:\mathbf{L}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}$ are defined the $L^{2}$ projection operators [32, 33] by
+$\left(v-I_{h}v,w_{h}\right)=0,~{}\forall w_{h}\in S_{h}^{r},$ (2.13)
+$\left(\mathbf{v}-\bm{I}_{h}\mathbf{v},\mathbf{w}_{h}\right)=0,~{}\forall\mathbf{w}_{h}\in\mathbf{X}_{h}^{r}.$ (2.14)
+It is known to all that the $L^{2}$ projection satisfies the following estimates,
+$\left\|v-I_{h}v\right\|+h\left\|\nabla\left(v-I_{h}v\right)\right\|\leq Ch^{r+1}\|v\|_{H^{r+1}},$ (2.15)
+$\left\|\mathbf{v}-\bm{I}_{h}\mathbf{v}\right\|+h\left\|\nabla\left(\mathbf{v}-\bm{I}_{h}\mathbf{v}\right)\right\|\leq Ch^{r+1}\|\mathbf{v}\|_{H^{r+1}}.$ (2.16)
+For $\left(\mathbf{v}_{h},q_{h}\right)\in\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$, the Stokes projection [34] $P_{h}$ or $\bm{P}_{h}:\mathbf{H}_{0}^{1}(\Omega)\times L_{0}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$ are defined by
+$\left(\nabla\left(\mathbf{u}-\bm{P}_{h}(\mathbf{u},p)\right),\nabla\mathbf{v}_{h}\right)-\left(p-P_{h}(\mathbf{u},p),\nabla\cdot\mathbf{v}_{h}\right)=0,$ (2.17)
+$\left(\nabla\cdot\left(\mathbf{u}-\bm{P}_{h}(\mathbf{u},p)\right),q_{h}\right)=0.$ (2.18)
+To be the simplicity of notations, we denote $P_{h}p=P_{h}(\mathbf{u},p)$ and $\bm{P}_{h}\mathbf{u}=\bm{P}_{h}(\mathbf{u},p)$.
+Lemma 2.1 ([35, 26]).
+Under the assumption of regularity (3.1), it holds that
+$\|\mathbf{u}-\bm{P}_{h}\mathbf{u}\|+h\left(\|\nabla\left(\mathbf{u}-\bm{P}_{h}\mathbf{u}\right)\|+\|p-P_{h}p\|\right)\leq Ch^{r+1}\left(\|\mathbf{u}\|_{H^{r+1}}+\|p\|_{H^{r}}\right),$ (2.19)
+and
+$\|\delta_{\tau}\left(\mathbf{u}^{n}-\bm{P}_{h}\mathbf{u}^{n}\right)\|\leq$ $~{}Ch^{r+1}\left(\|\delta_{\tau}\mathbf{u}^{n}\|_{H^{r+1}}+\|\delta_{\tau}p^{n}\|_{H^{r}}\right).$ (2.20)
+where $C$ is the constant the only depends on $\Omega$.
+The proof of this lemma is similar to Lemma 3.2 of [26]. Here we omit the detailed procedure and refer the interested reader to [26].
+Lemma 2.2 ([36]).
+The Stokes projection $\bm{P}_{h}(\mathbf{u},p)$ is $\mathbf{H}^{1}$ stable in the case such that
+$\|\bm{P}_{h}(\mathbf{u},p)\|_{H^{1}}\leq C\left(\|\mathbf{u}\|_{H^{2}}+\|p\|_{H^{1}}\right).$ (2.21)
+Next we introduce the discrete Laplacian operator [37] $\Delta_{h}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ such that
+$\left(-\Delta_{h}v_{h},\xi_{h}\right)=\left(\nabla v_{h},\nabla\xi_{h}\right),$ (2.22)
+$\left(\nabla(-\Delta_{h}^{-1})v_{h},\nabla\xi_{h}\right)=\left(v_{h},\xi_{h}\right).$ (2.23)
+By defining the discrete Stokes operator $A_{h}=-\bm{I}_{h}\Delta_{h}$ [27], it holds
+$\|v_{h}\|_{H^{2}}=\|A_{h}v_{h}\|,~{}\|v_{h}\|_{H^{-1}}=\|A_{h}^{-1/2}v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r},$ (2.24)
+$\|A_{h}^{1/2}v_{h}\|=\|\nabla v_{h}\|,~{}\|\nabla A_{h}^{-1/2}v_{h}\|=\|v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r}.$ (2.25)
+If $v_{h}$ is a constant, we denote $(-\Delta_{h})^{\frac{1}{2}}v_{h}=0$ and $(-\Delta_{h})^{-\frac{1}{2}}v_{h}=0$.
+Lemma 2.3 ([26]).
+For $v_{h}\in\mathring{S}_{h}^{r-1}$, and the operators $\left(-\Delta_{h}\right)^{1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ and $\left(-\Delta_{h}\right)^{-1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$, we have the estimates
+$\|v_{h}\|_{H^{1}}\leq C\|\left(-\Delta_{h}\right)^{1/2}v_{h}\|\leq c\|v_{h}\|_{H^{1}},$ (2.26)
+$\|v_{h}\|_{H^{-1}}\leq C\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|\leq c\|v_{h}\|_{H^{-1}},$ (2.27)
+and
+$\|v_{h}\|^{2}\leq\varepsilon\|\nabla v_{h}\|^{2}+C\varepsilon^{-1}\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|^{2},$ (2.28)
+where $\varepsilon$ is an arbitrary positive constant independent of $h$.
+Finally, we note the well-known inverse inequalities [26]
+$\|v_{h}\|_{L^{4}}\leq Ch^{-1/2}\|v_{h}\|,\quad\forall v_{h}\in S_{h}^{r},$ (2.29)
+$\|v_{h}\|_{L^{\infty}}\leq C\sqrt{\ln(1/h)}\|v_{h}\|_{H^{1}},\quad\forall v_{h}\in S_{h}^{r}.$ (2.30)
+Lemma 2.4 ([38, 39]).
+Suppose $g\in H^{1}(\Omega)$, and $v\in\mathring{S}_{h}^{r-1}$,
+$|(g,v)|\leq C\|\nabla g\|\|v\|_{H^{-1}},$ (2.31)
+and for all $g\in L_{0}^{2}(\Omega)$, we have
+$\|g\|_{H^{-1}}\leq C\|g\|.$ (2.32)
+where $C$ is positive constant independent of $\tau$ and $h$.
+The proof of the lemma 2.4 is similar to Lemma 3.2 of [38], here we omit it. The following inequalities hold [40]
+$\|\phi\|_{L^{r}}\leq C\|\phi\|_{H^{1}},~{}\forall\phi\in H^{1}(\Omega),~{}r\in[2,\infty],$ (2.33)
+$\|\phi\|_{L^{\infty}}\leq C\|\phi\|^{\frac{1}{2}}\left(\|\phi\|^{2}+\|\Delta\phi\|^{2}\right)^{\frac{1}{2}},\forall\phi\in H^{2}(\Omega),$ (2.34)
+$\|\bm{v}\|_{L^{r}}\leq C\|\nabla\bm{v}\|^{2},~{}(2\leq r\leq 6),\quad\|\bm{v}\|_{L^{4}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|\nabla\bm{v}\|^{\frac{1}{2}},~{}\forall\bm{v}\in\mathbf{H}_{0}^{1}(\Omega),$ (2.35)
+$\|\bm{v}\|_{H^{2}}\leq C\|A\bm{v}\|^{2},\quad\|\bm{v}\|_{L^{\infty}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|A\bm{v}\|^{\frac{1}{2}},\forall\bm{v}\in\mathbf{H}^{2}(\Omega)\cap\mathbf{H}_{0}^{1}(\Omega),$ (2.36)
+where $A$ is the continuous Stokes operator [34].
 We will frequently use the following discrete version of the Grönwall lemma [41]:
@@
 $\left(\nabla\cdot\mathbf{u}_{h}^{n+1},q_{h}\right)=0,~{}\mathbf{u}_{h}^{n+1}\cdot\mathbf{n}|_{\partial\Omega}=0,$ (3.12)
-hold for all $\left(w_{h},\varphi_{h},\mathbf{v}_{h},q_{h}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$, and $n=0,1,2,\cdots,N-1$, where $\phi_{h}^{0}=R_{h}\phi^{0},~{}\mathbf{u}_{h}^{0}=\mathbf{I}_{h}\mathbf{u}^{0}.$ For simplicity of description, we perform the fully discretization to prove that our numerical scheme satisfies energy stabilization.
+hold for all $\left(w_{h},\varphi_{h},\mathbf{v}_{h},q_{h}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$, and $n=0,1,2,\cdots,N-1$, where $\phi_{h}^{0}=R_{h}\phi^{0},~{}\mathbf{u}_{h}^{0}=\bm{I}_{h}\mathbf{u}^{0}.$ For simplicity of description, we perform the fully discretization to prove that our numerical scheme satisfies energy stabilization.
 For the convenience of the following theoretical analysis, we set the parameters $\lambda,M,\nu$ to 1.
-3.2 Energy stability
-For the above scheme, we can establish the theorem of energy stabilization as follows:
+3.2 Unique solvability and energy stability
+For the fully discrete scheme (3.6)-(3.12), we assume that the initial solutions satisfly the following boundedness:
+$\left(F(\phi_{h}^{0}),1\right)+\|\phi_{h}^{0}\|_{H^{1}}+\|\mathbf{u}_{h}^{0}\|\leq C,$ (3.13)
+where $C$ is a constant independent of mesh size $h$.
+In this subsection, we establish the existence of solutions for the fully discrete scheme (3.6)-(3.12).
+This is achieved by applying the Leray–Schauder fixed point theorem, and we also prove the uniqueness of the numerical solutions.
+Given that the scheme (3.6)-(3.12) is implemented in a stepwise, decoupled manner, the existence of a solution is demonstrated incrementally.
+For the Step 2 and Step 3 of the discrete scheme, the proofs of existence and uniqueness are standard results; thus, we omit the detailed discussions here.
+Instead, our focus is on proving the existence and uniqueness of the solution for Step 1, specifically for the fully discrete scheme (3.6)-(3.7) corresponding to the CH equation.
+Lemma 3.1 ([42]).
+Let $\mathcal{G}$ denote a compact mapping from a Banach space $B$ into itself, and assume that there exists a constant $C$ such that $\|x\|<C$ for all $x\in B$ and $\alpha\in[0,1]$ satisfying $x=\alpha\mathcal{G}x$.
+Thus, $\mathcal{G}$ has a fixed point.
+Lemma 3.2.
+Suppose the initial solutions boundedness (3.13) holds.
+For given $\tau,h>0$, there exists a solution $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)$ to the fully discrete scheme (3.6)-(3.12).
+Lemma 3.3.
+For $\tau,h>0$ and $\tau\leq C\alpha/{\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the fully discrete scheme (3.6)-(3.12) is unique solvability.
+The proofs of Lemmas 3.2 and 3.3 are given in Appendix 6.1 and 6.2.
+Remark 3.1.
+According to the assumption $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the above Lemma (3.3) holds.
+Therefore, we can derive the boundedness of the nth-time-layer velocity as follows:
+$\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}\leq C\|\mathbf{u}\|_{L^{\infty}(0,T;\mathbf{L}^{\infty}(\Omega))}\leq C,$ (3.14)
+where $C$ is a positive constant independent of $\alpha$, $\tau$, $h$, and $n$.
+Later in this paper, we will analyze the reasonableness of this condition $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$.
+We explain later in Remark 4.1 the feasibility of showing that the inequality holds. For the above scheme, we can establish the theorem of energy stabilization as follows:
 Theorem 3.1.
 Let $\phi_{h}^{n}\in S_{h}^{r}$ and $\mathbf{u}_{h}^{n}\in\mathbf{X}_{h}^{r}$ and the initial values $|\phi_{h}^{0}|\leq C_{1}$ and $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})\leq C_{0}$, i.e., the initial energy is bounded.
-For all time partitions $\tau$ and mesh size $h$ such that $\tau\leq\min\{Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\}$, the decoupled scheme (3.6)-(3.12) is uniquely solvable and satisfies discrete energy law as follows:
+For all time partitions $\tau$ and mesh size $h$ such that $\tau\leq C\alpha/{\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$ and the second step satisfies $\tau\leq Ch/\|\phi_{h}^{n+1}\|_{L^{4}}^{2}$, the decoupled discrete scheme (3.6)-(3.12) is uniquely solvable and satisfies discrete energy law as follows:
 $\tilde{E}_{h}^{n+1}-\tilde{E}_{h}^{n}\leq-M\tau\|\nabla\mu_{h}^{n+1}\|^{2}-2\nu\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2},$ (3.15)
@@
 Combining the (3.17), (3.18) with (3.22), we get
-$\lambda\left(\|\nabla\phi_{h}^{n+1}\|^{2}\right.$ $\left.-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.21)
-$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$
+$\lambda$ $\left(\|\nabla\phi_{h}^{n+1}\|^{2}-\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}\right)$ (3.23)
+$+\frac{2\lambda\tau}{\epsilon^{2}}\left(\frac{1}{4}\delta_{\tau}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\frac{\tau}{4}\left(\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}\right.\right.$
+$\left.\left.\qquad+2\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+2\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right)\right)$
 $+\|\mathbf{u}_{h}^{n+1}\|^{2}-\|\mathbf{u}_{h}^{n}\|^{2}+\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+\tau^{2}\left(\|\nabla p_{h}^{n+1}\|^{2}-\|\nabla p_{h}^{n}\|^{2}\right)$
 $=$ $-2M\tau\|\nabla\mu_{h}^{n+1}\|^{2}-2\nu\tau\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+2\tau\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right).$
-Recalling inverse inequalities and choosing $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, we have
-$2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\bigg{|}\leq$ $~{}2\tau\|\phi_{h}^{n+1}\|_{L^{4}}\|\nabla\mu_{h}^{n+1}\|\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}$ (3.22)
-$\leq$ $~{}C\tau\|\phi_{h}^{n+1}-\phi_{h}^{n}+\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2}$
-$\leq$ $~{}\frac{1}{2}\|\nabla\phi_{h}^{n+1}-\nabla\phi_{h}^{n}\|^{2}+\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2},$
+According to (3.3), we obtain $\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$. Recalling inverse inequality and choosing $\tau\leq Ch$, we have
+$2\tau\bigg{|}\left(\mu_{h}^{n+1}\nabla\phi_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\right)\bigg{|}$ (3.24)
+$\leq$ $~{}C\tau\|\nabla\mu_{h}^{n+1}\|\|\phi_{h}^{n+1}\|_{L^{4}}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|_{L^{4}}$
+$\leq$ $~{}\frac{C_{2}\tau}{h}\|\phi_{h}^{n+1}\|_{L^{4}}^{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+\tau\|\nabla\mu_{h}^{n+1}\|^{2}$
+$\leq$ $~{}\frac{1}{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}+M\tau\|\nabla\mu_{h}^{n+1}\|^{2},$
 where $C$ and $C_{2}$ are constant that does not depend on $\tau$ and $h$. Combining (3.24) with (3.23), and summing up form $n=0$ to $m$, we have
-$\tilde{E}_{h}^{m+1}-\tilde{E}_{h}^{0}\leq-\tau\sum_{n=0}^{m}\left(2\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}\right)-\sum_{n=0}^{m}\left(1-\frac{C_{2}\tau}{h}\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}.$ (3.23)
-In order to the energy to satisfy the law of decay when $m=0$, the following conditions need to be satisfied
-$\tau\leq\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}}.$ (3.24)
-Since $\|\phi_{h}^{0}\|_{L^{4}}\leq C\|\phi_{h}^{0}\|$ is bounded, the $\tau$ of the above equation (3.24) can be taken to a valid value such that
-$\tilde{E}_{h}^{1}-\tilde{E}_{h}^{0}\leq 0.$ (3.25)
-So we can obtain that $\|\phi_{h}^{1}\|$ is bounded. When $m=1$, we have
-$\tau\leq\min\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}}\bigg{\}},$ (3.26)
-then we can dervie that $\|\phi_{h}^{2}\|\leq C$. By analogy, we can obatin the following result
-$\tau\leq\min_{0\leq n\leq m}\bigg{\{}\frac{Ch}{\|\phi_{h}^{0}\|_{L^{4}}^{2}},~{}\frac{Ch}{\|\phi_{h}^{1}\|_{L^{4}}^{2}},~{}\cdots,~{}\frac{Ch}{\|\phi_{h}^{n}\|_{L^{4}}^{2}}\bigg{\}}.$ (3.27)
-Thus, we can always obtain a positive number $C$ such that $\|\phi_{h}^{n}\|_{L4}\leq C$ is bounded, and derive the desired result (3.13). ∎
+$\tilde{E}_{h}^{m+1}-\tilde{E}_{h}^{0}\leq-\tau\sum_{n=0}^{m}\left(2\|\nabla\tilde{\mathbf{u}}_{h}^{n+1}\|^{2}+\|\nabla\mu_{h}^{n+1}\|^{2}\right)-\sum_{n=0}^{m}\frac{1}{2}\|\tilde{\mathbf{u}}_{h}^{n+1}-\mathbf{u}_{h}^{n}\|^{2}.$ (3.25)
+Thus, we can always obtain a positive number $C$ that does not mesh size $h$ and time step $\tau$ such that $\|\phi_{h}^{n+1}\|_{L^{4}}\leq\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$ is bounded, and derive the desired result (3.15).
+∎
 Remark 3.2.
@@
 Let $\left(\phi_{h}^{n+1},\mu_{h}^{n+1},\tilde{\mathbf{u}}_{h}^{n+1},\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)\in\mathbf{\mathcal{X}}_{h}^{r}$ be the unique solution of (3.6)-(3.12).
-Suppose that $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})<C_{0}$, for any $h,\tau>0$ and $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$, the following boundedness hold that
+Suppose that $E(\phi_{h}^{0},\mathbf{u}_{h}^{0})<C_{0}$, for any $h,\tau>0$ and the time step $\tau$ and mesh size $h$ satisfy the conditions of Theorem 3.1, the following boundednesses hold
 $\max_{0\leq n\leq N-1}\left(\|\mathbf{u}_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}+\frac{\tau}{2\epsilon^{2}}\|\left(\phi_{h}^{n+1}\right)^{2}-1\|^{2}+\tau^{2}\|\nabla p_{h}^{n+1}\|^{2}\right)$ $\leq C_{0},$ (3.26)
@@
 $\frac{\tau^{2}}{\epsilon^{2}}\sum_{n=0}^{N-1}\left[\frac{1}{2}\|\delta_{\tau}(\phi_{h}^{n+1})^{2}\|^{2}+\|\phi_{h}^{n+1}\delta_{\tau}\phi_{h}^{n+1}\|^{2}+\|\delta_{\tau}\phi_{h}^{n+1}\|^{2}\right]$ $\leq C_{0},$ (3.29)
-where $C_{0}$ is a positive constant dependent on the initial value of $\phi^{0}$ and $\mathbf{u}^{0}$.
+where $C_{0}$ is a positive constant dependent on the initial values of $\phi^{0}$ and $\mathbf{u}^{0}$.
 The proof of this lemma can be derived from the result of Theorem 3.1, the detailed steps of which we omit.
@@
 $\|\phi_{h}^{m+1}\|^{2}\leq\|\phi_{h}^{0}\|^{2}+\tau\sum_{n=0}^{m}\left(\|\nabla\mu_{h}^{n+1}\|^{2}+2\|\nabla\phi_{h}^{n+1}\|^{2}+C\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\phi_{h}^{n+1}\|^{2}+\|\nabla\phi_{h}^{n+1}\|^{2}\right)\right).$ (3.35)
-Using the lemma 2.1 and 3.1, we have
+Using the lemma 2.5 and 3.4, we have
 $\|\phi_{h}^{m+1}\|^{2}\leq\left(\|\phi_{h}^{0}\|^{2}+c_{1}\left(1+4\tau+c_{2}\right)\right)\exp(c)\leq C\tau.$ (3.36)
@@
 $\|\Delta_{h}\phi_{h}^{m+1}\|^{2}+\tau\sum_{n=0}^{m}\|\Delta_{h}\mu_{h}^{n+1}\|^{2}\leq\|\Delta_{h}\phi_{h}^{0}\|^{2}+\tau\sum_{n=0}^{m}\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\nabla\phi_{h}^{n+1}\|^{2}+\|\Delta_{h}\phi_{h}^{n+1}\|^{2}\right).$ (3.44)
-Using the lemma 2.1 and 3.1, we obtain the result of this lemma. ∎
+Using the lemma 2.5 and 3.4, we obtain the result of this lemma. ∎
 Remark 3.3.
-We note that the bound $\tau\leq Ch/\left(2\|\phi_{h}^{n}\|_{L^{4}}^{2}\right)$ in Lemma 3.1 and 3.2 is attainable, since from (4.26) and Lemma 3.2 we obtain $\|\phi_{h}^{n}\|_{L^{4}}\leq C\|\phi_{h}^{n}\|_{H^{1}}\leq C$.
+We note that the bound $\tau\leq Ch/\left(2\|\phi_{h}^{n+1}\|_{L^{4}}^{2}\right)$ in Lemma 3.4 and 3.5 is attainable, since from (2.33) and Lemma 3.5 and (6.3), we obtain $\|\phi_{h}^{n+1}\|_{L^{4}}\leq C\|\phi_{h}^{n+1}\|_{H^{1}}\leq C$.
 More similar details about energy stabilization under this discrete scheme can be found in reference [4].
 4 Error analysis
-In this section, we establish the error analysis for the scheme (3.6)-(3.12). Firstly, we denote $R_{h}:H^{1}(\Omega)\rightarrow S_{h}^{r}$ as the classic Ritz projection [30],
-$\left(\nabla\left(\psi-R_{h}\psi\right),\nabla\varphi_{h}\right)=0,\quad\forall~{}\psi\in H^{1}(\Omega),~{}\varphi_{h}\in S_{h}^{r}.$ (4.1)
-Now, for $s\in[2,\infty]$ and $n=1,2,\cdots,N$, from the finite element approximation theory [31], it holds that
-$\|\psi-R_{h}\psi\|_{L^{s}}+h\|\psi-R_{h}\psi\|_{W^{1,s}}\leq Ch^{r+1}\|\psi\|_{W^{r+1,s}},$ (4.2)
-$\|\psi-R_{h}\psi\|_{H^{-1}}\leq Ch^{r+1}\|\psi\|_{H^{r+1}},$ (4.3)
-$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|+h\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (4.4)
-$\|\delta_{\tau}\left(\psi^{n}-R_{h}\psi^{n}\right)\|_{H^{-1}}\leq Ch^{r+1}\|\delta_{\tau}\psi^{n}\|_{H^{r+1}},$ (4.5)
-where $\delta_{\tau}$ is defined in (2.7).
-Let $I_{h}:L^{2}(\Omega)\rightarrow S_{h}^{r}$ and $\mathbf{I}_{h}:\mathbf{L}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}$ are defined the $L^{2}$ projection operators [32, 33] by
-$\left(v-I_{h}v,w_{h}\right)=0,~{}\forall w_{h}\in S_{h}^{r},$ (4.6)
-$\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v},\mathbf{w}_{h}\right)=0,~{}\forall\mathbf{w}_{h}\in\mathbf{X}_{h}^{r}.$ (4.7)
-It is known to all that the $L^{2}$ projection satisfies the following estimates,
-$\left\|v-I_{h}v\right\|+h\left\|\nabla\left(v-I_{h}v\right)\right\|\leq Ch^{r+1}\|v\|_{H^{r+1}},$ (4.8)
-$\left\|\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right\|+h\left\|\nabla\left(\mathbf{v}-\mathbf{I}_{h}\mathbf{v}\right)\right\|\leq Ch^{r+1}\|\mathbf{v}\|_{H^{r+1}}.$ (4.9)
-For $\left(\mathbf{v}_{h},q_{h}\right)\in\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$, the Stokes projection [34] $P_{h}$ or $\mathbf{P}_{h}:\mathbf{H}_{0}^{1}(\Omega)\times L_{0}^{2}(\Omega)\rightarrow\mathbf{X}_{h}^{r}\times\mathring{S}_{h}^{r-1}$ are defined by
-$\left(\nabla\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),\nabla\mathbf{v}_{h}\right)-\left(p-P_{h}(\mathbf{u},p),\nabla\cdot\mathbf{v}_{h}\right)=0,$ (4.10)
-$\left(\nabla\cdot\left(\mathbf{u}-\mathbf{P}_{h}(\mathbf{u},p)\right),q_{h}\right)=0.$ (4.11)
-To be the simplicity of notations, we denote $P_{h}p=P_{h}(\mathbf{u},p)$ and $\mathbf{P}_{h}\mathbf{u}=\mathbf{P}_{h}(\mathbf{u},p)$.
-Lemma 4.1 ([35, 25]).
-Under the assumption of regularity (3.1), it holds that
-$\|\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\|+h\left(\|\nabla\left(\mathbf{u}-\mathbf{P}_{h}\mathbf{u}\right)\|+\|p-P_{h}p\|\right)\leq Ch^{r+1}\left(\|\mathbf{u}\|_{H^{r+1}}+\|p\|_{H^{r}}\right),$ (4.12)
-and
-$\|\delta_{\tau}\left(\mathbf{u}^{n}-\mathbf{P}_{h}\mathbf{u}^{n}\right)\|\leq$ $~{}Ch^{r+1}\left(\|\delta_{\tau}\mathbf{u}^{n}\|_{H^{r+1}}+\|\delta_{\tau}p^{n}\|\right.$ (4.13)
-$~{}\left.+\|\delta_{\tau}\phi^{n}\|_{H^{r+1}}\|\mu^{n}\|_{H^{2}}+\|\phi^{n-1}\|_{H^{r+1}}\|\delta_{\tau}\mu^{n}\|_{H^{2}}\right).$
-where $C$ is the constant the only depends on $\Omega$.
-The proof of this lemma is similar to Lemma 3.2 of [25]. Here we omit the detailed procedure and refer the interested reader to [25].
-Lemma 4.2 ([36]).
-The Stokes projection $\mathbf{P}_{h}(\mathbf{u},p)$ is $\mathbf{H}^{1}$ stable in the case such that
-$\|\mathbf{P}_{h}(\mathbf{u},p)\|_{H^{1}}\leq C\left(\|\mathbf{u}\|_{H^{2}}+\|p\|_{H^{1}}\right).$ (4.14)
-Next we introduce the discrete Laplacian operator [37] $\Delta_{h}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ such that
-$\left(-\Delta_{h}v_{h},\xi_{h}\right)=\left(\nabla v_{h},\nabla\xi_{h}\right),$ (4.15)
-$\left(\nabla(-\Delta_{h}^{-1})v_{h},\nabla\xi_{h}\right)=\left(v_{h},\xi_{h}\right).$ (4.16)
-By defining the discrete Stokes operator $A_{h}=-\mathbf{I}_{h}\Delta_{h}$ [26], it holds
-$\|v_{h}\|_{H^{2}}=\|A_{h}v_{h}\|,~{}\|v_{h}\|_{H^{-1}}=\|A_{h}^{-1/2}v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r},$ (4.17)
-$\|A_{h}^{1/2}v_{h}\|=\|\nabla v_{h}\|,~{}\|\nabla A_{h}^{-1/2}v_{h}\|=\|v_{h}\|,~{}v_{h}\in\mathbf{X}_{h}^{r}.$ (4.18)
-If $v_{h}$ is a constant, we denote $(-\Delta_{h})^{\frac{1}{2}}v_{h}=0$ and $(-\Delta_{h})^{-\frac{1}{2}}v_{h}=0$.
-Lemma 4.3 ([25]).
-For $v_{h}\in\mathring{S}_{h}^{r-1}$, and the operators $\left(-\Delta_{h}\right)^{1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$ and $\left(-\Delta_{h}\right)^{-1/2}:\mathring{S}_{h}^{r-1}\rightarrow\mathring{S}_{h}^{r-1}$, we have the estimates
-$\|v_{h}\|_{H^{1}}\leq C\|\left(-\Delta_{h}\right)^{1/2}v_{h}\|\leq c\|v_{h}\|_{H^{1}},$ (4.19)
-$\|v_{h}\|_{H^{-1}}\leq C\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|\leq c\|v_{h}\|_{H^{-1}},$ (4.20)
-and
-$\|v_{h}\|^{2}\leq\varepsilon\|\nabla v_{h}\|^{2}+C\varepsilon^{-1}\|\left(-\Delta_{h}\right)^{-1/2}v_{h}\|^{2},$ (4.21)
-where $\varepsilon$ is an arbitrary positive constant independent of $h$.
-Finally, we note the well-known inverse inequalities [25]
-$\|v_{h}\|_{L^{4}}\leq Ch^{-1/2}\|v_{h}\|,\quad\forall v_{h}\in S_{h}^{r},$ (4.22)
-$\|v_{h}\|_{L^{\infty}}\leq C\sqrt{\ln(1/h)}\|v_{h}\|_{H^{1}},\quad\forall v_{h}\in S_{h}^{r}.$ (4.23)
-Lemma 4.4 ([38, 29]).
-Suppose $g\in H^{1}(\Omega)$, and $v\in\mathring{S}_{h}^{r-1}$,
-$|(g,v)|\leq C\|\nabla g\|\|v\|_{H^{-1}},$ (4.24)
-and for all $g\in L_{0}^{2}(\Omega)$, we have
-$\|g\|_{H^{-1}}\leq C\|g\|.$ (4.25)
-where $C$ is positive constant independent of $\tau$ and $h$.
-The proof of the lemma 4.4 is similar to Lemma 3.2 of [38], here we omit it. The following inequalities hold [39]
-$\|\phi\|_{L^{r}}\leq C\|\phi\|_{H^{1}},~{}\forall\phi\in H^{1}(\Omega),~{}r\in[2,\infty],$ (4.26)
-$\|\phi\|_{L^{\infty}}\leq C\|\phi\|^{\frac{1}{2}}\left(\|\phi\|^{2}+\|\Delta\phi\|^{2}\right)^{\frac{1}{2}},\forall\phi\in H^{2}(\Omega),$ (4.27)
-$\|\bm{v}\|_{L^{r}}\leq C\|\nabla\bm{v}\|^{2},~{}(2\leq r\leq 6),\quad\|\bm{v}\|_{L^{4}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|\nabla\bm{v}\|^{\frac{1}{2}},~{}\forall\bm{v}\in\mathbf{H}_{0}^{1}(\Omega),$ (4.28)
-$\|\bm{v}\|_{H^{2}}\leq C\|A\bm{v}\|^{2},\quad\|\bm{v}\|_{L^{\infty}}\leq C\|\bm{v}\|^{\frac{1}{2}}\|A\bm{v}\|^{\frac{1}{2}},\forall\bm{v}\in\mathbf{H}^{2}(\Omega)\cap\mathbf{H}_{0}^{1}(\Omega),$ (4.29)
-where $A$ is the continuous Stokes operator [34].
-And then, we denote
+In this section, we establish the error analysis for the scheme (3.6)-(3.12). We denote
 $e_{\phi}^{n+1}=\phi^{n+1}-\phi_{h}^{n+1}=\phi^{n+1}-R_{h}\phi^{n+1}+R_{h}\phi^{n+1}-\phi_{h}^{n+1}=\Phi_{\phi h}^{n+1}+\Theta_{\phi h}^{n+1},$
@@
 and
-$e_{\tilde{\mathbf{u}}}^{n+1}=\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\tilde{\mathbf{u}}h}^{n+1},$
-$e_{\mathbf{u}}^{n+1}=\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\mathbf{u}^{n+1}-\mathbf{P}_{h}\mathbf{u}^{n+1}+\mathbf{P}_{h}\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\mathbf{u}h}^{n+1},$
+$e_{\tilde{\mathbf{u}}}^{n+1}=\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\mathbf{u}^{n+1}-\bm{P}_{h}\mathbf{u}^{n+1}+\bm{P}_{h}\mathbf{u}^{n+1}-\tilde{\mathbf{u}}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\tilde{\mathbf{u}}h}^{n+1},$
+$e_{\mathbf{u}}^{n+1}=\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\mathbf{u}^{n+1}-\bm{P}_{h}\mathbf{u}^{n+1}+\bm{P}_{h}\mathbf{u}^{n+1}-\mathbf{u}_{h}^{n+1}=\Phi_{\mathbf{u}h}^{n+1}+\Theta_{\mathbf{u}h}^{n+1},$
 $e_{p}^{n+1}=p^{n+1}-p_{h}^{n+1}=p^{n+1}-P_{h}p^{n+1}+P_{h}p^{n+1}-p_{h}^{n+1}=\Phi_{ph}^{n+1}+\Theta_{ph}^{n+1}.$
@@
 where the truncation errors $R_{1}^{n+1}$, $R_{2}^{n+1}$ and $R_{3}^{n+1}$ is denoted by
-$R_{1}^{n+1}=\delta_{\tau}R_{h}\phi^{n+1}-\partial_{t}\phi^{n+1},R_{2}^{n+1}=\delta_{\tau}\mathbf{P}_{h}\mathbf{u}^{n+1}-\partial_{t}\mathbf{u}^{n+1},R_{3}^{n+1}=\nabla\left(P_{h}p^{n+1}-P_{h}p^{n}\right),$ (4.35)
+$R_{1}^{n+1}=\delta_{\tau}R_{h}\phi^{n+1}-\partial_{t}\phi^{n+1},R_{2}^{n+1}=\delta_{\tau}\bm{P}_{h}\mathbf{u}^{n+1}-\partial_{t}\mathbf{u}^{n+1},R_{3}^{n+1}=\nabla\left(P_{h}p^{n+1}-P_{h}p^{n}\right),$ (4.6)
 and the $\delta_{\tau}$ in time scheme is defined by (2.7).
@@
 Under the assumption of regularity (3.1), for any $\tau\geq 0$ and $n=0,1,2,\cdots,N-1$, the truncation errors satisfy
-$\|R_{1}^{n+1}\|+\|R_{2}^{n+1}\|+\|R_{3}^{n+1}\|+\|R_{4}^{n+1}\|\leq C\tau.$ (4.40)
+$\tau\sum_{n=0}^{n}\left(\|R_{1}^{n+1}\|^{2}+\|R_{2}^{n+1}\|^{2}+\|R_{3}^{n+1}\|^{2}+\|R_{4}^{n+1}\|^{2}\right)\leq C\tau^{2}.$ (4.11)
 The detailed proof of this Lemma can refer to [38, 39].
@@
 $\delta_{\tau}\Theta_{\phi h}^{n+1}=\Delta_{h}e_{\mu}^{n+1}-\delta_{\tau}\Phi_{\phi h}^{n+1}-\left(\mathbf{u}^{n+1}\cdot\nabla\phi^{n+1}-\mathbf{u}_{h}^{n}\cdot\nabla\phi_{h}^{n+1}\right)+R_{1}^{n+1}.$ (4.18)
-Using the (4.26), (4.28), Lemma 4.1 and 4.2, we can obtain
+Using the (2.33), (2.35), Lemma 2.1 and 2.2, we can obtain
 $\|\delta_{\tau}\Theta_{\phi h}^{n+1}\|_{H^{-1}}\leq C\|\nabla e_{\mu}^{n+1}\|+\|R_{1}^{n+1}\|$ (4.19)
@@
 $\left.\qquad+\|\Theta_{\mathbf{u}h}^{n+1}\|^{2}+\|\Phi_{\phi h}^{n+1}\|_{H^{1}}^{2}+\|\Theta_{\phi h}^{n+1}\|_{H^{1}}^{2}\right).$
-Using the (4.2), and Lemma 4.1, 3.1, 4.5, and 2.1, we can obtain the result of this Lemma. ∎
+Using the (2.9), and Lemma 2.1, 3.4, 4.1, and 2.5, we can obtain the result of this Lemma. ∎
 Lemma 4.4.
@@
 $+C\tau^{3}\sum_{n=0}^{m}\|\phi^{n+1}\|_{H^{2}}^{2}\left(\|\nabla\Theta_{ph}^{n+1}\|^{2}+\|\nabla\Theta_{ph}^{n}\|^{2}+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}^{2}\left(\|A_{h}\mathbf{u}^{n+1}\|^{2}+\|A_{h}\mathbf{u}^{n}\|^{2}\right)\right).$
-According to the Lemma 4.1 and 4.2, we can derive
+According to the Lemma 2.1 and 2.2, we can derive
 […]
@@
 […]
-Finally, using the Lemma 4.1, 3.1, 3.2, 4.5, the inequalities (4.2)-(4.5) and the discrete Grönwall’s lemma 2.1, we can ontain the proof of this lemma. ∎
+Finally, using the Lemma 2.1, 3.4, 3.5, 4.1, the inequalities (2.9)-(2.12) and the discrete Grönwall’s lemma 2.5, we can ontain the proof of this lemma. ∎
 Therefore, we can obtain the following theorem based on the above lemmas.
@@
 $+C\tau\sum_{n=0}^{m}\left(\left(\|\nabla\mathbf{u}^{n+1}\|^{2}+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\right)\left(\|\Theta_{{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)+\|\nabla\mathbf{u}_{h}^{n}\|^{2}\left(\|\Theta_{\tilde{\mathbf{u}}h}^{n+1}\|^{2}+\|\Phi_{\mathbf{u}h}^{n+1}\|^{2}\right)\right).$
-According to the triangle inequality, the inequality (4.3), the Lemma 4.1, 3.1 and 3.2, this completes the proof of Lemma 4.9. ∎
+According to the triangle inequality, the inequality (2.10), the Lemma 2.1, 3.4 and 3.5, this completes the proof of Lemma 4.5. ∎
+Remark 4.1.
+Finally, we aim to confirm that, as a consequence of $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the $a~{}priori$ assumption (6.7) remains valid for all sufficiently small values of $h$.
+To achieve this, we will show that, for sufficiently small $h$, the following requirement is satisfied.
+We proceed to prove (6.7) using mathematical induction.
+Specifically, for $h\leq h_{2}$, (6.7) holds when $n=0$, given that $\mathbf{u}_{h}^{0}=\bm{P}_{h}\mathbf{u}^{0}$.
+Assuming that (6.7) holds for $n\leq m$, we will demonstrate that this assumption also holds for $n=m+1$.
+Indeed, if (6.7) holds for $n\leq m$, then (6.6) holds for $n=m$.
+By using the inverse inequality, for $h\leq h_{2}$, we have
+$\left\|\mathbf{u}_{h}^{m+1}\right\|_{L^{\infty}}\leq$ $~{}\left\|\mathbf{u}_{h}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}+\left\|\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}$ (4.55)
+$\leq$ $~{}C_{1}h^{-d/2}\left\|\mathbf{u}_{h}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|+\left\|\bm{P}_{h}\mathbf{u}^{m+1}\right\|_{L^{\infty}}$
+$\leq$ $~{}C_{1}h^{-d/2}\left\|\mathbf{u}^{m+1}-\mathbf{u}_{h}^{m+1}\right\|+Ch^{-d/2}\left\|\mathbf{u}^{m+1}-\bm{P}_{h}\mathbf{u}^{m+1}\right\|$
+$\quad+C\|\mathbf{u}\|_{L^{\infty}\left(0,T;\mathbf{L}^{\infty}(\Omega)\right)}$
+$\leq$ $~{}C\|\mathbf{u}\|_{L^{\infty}\left(0,T;\mathbf{L}^{\infty}(\Omega)\right)}.$
+Therefore, (6.7) also holds at $n=m+1$. The boundedness explanation for $\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}$ is complete.
 Lemma 4.6.
 Under the assumption of regularity (3.1), for all $m\geq 0$, there exists a positive constant that does not depend on $\tau$ and $h$ such that
-$\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}\leq C\tau\left(\tau^{2}+h^{2r}\right).$ (4.84)
+$\sum_{n=0}^{m}\|e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\|_{H^{-1}}^{2}\leq C\left(\tau^{2}+h^{2r}\right).$ (4.56)
 Proof.
@@
 $+\tau\left(R_{2}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right)+\tau\left(R_{3}^{n+1},A_{h}^{-1}(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n})\right).$
-According to (4.6) and (3.11), we can obtain $\mathbf{u}_{h}^{n+1}=\mathbf{I}_{h}\tilde{\mathbf{u}}_{h}^{n+1}$. Using the (4.15) and the above results, we have
+According to (2.13) and (3.11), we can obtain $\mathbf{u}_{h}^{n+1}=\bm{I}_{h}\tilde{\mathbf{u}}_{h}^{n+1}$. Using the (2.22) and the above results, we have
 $\tau\left(\Delta_{h}e_{\tilde{\mathbf{u}}}^{n+1},A_{h}^{-1}\left(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)\right)=$ $~{}-\left(e_{\tilde{\mathbf{u}}}^{n+1},-\tau\Delta_{h}A_{h}^{-1}\left(e_{\mathbf{u}}^{n+1}-e_{\mathbf{u}}^{n}\right)\right)$ (4.59)
@@
 $+\|\nabla\mathbf{u}^{n+1}\|\left(\|e_{\mathbf{u}}^{n}\|+\|\mathbf{u}_{t}\|_{L^{\infty}(0,T;\mathbf{H}^{1}(\Omega))}\right)+\|e_{\mu}^{n+1}\|_{H^{1}}\|\nabla\phi^{n+1}\|+\|\mu_{h}^{n+1}\|_{H^{1}}\|\nabla e_{\phi}^{n+1}\|.$
-Squaring both ends of the (4.93) and multiplying the result by $\tau$, summing over $n~{}(0\leq n\leq m)$, and using the Lemma 3.1, 3.2, 4.5, 4.10 and the Theorem 4.2, we can obtain the result of this lemma as follows
+Squaring both ends of the (4.65) and multiplying the result by $\tau$, summing over $n~{}(0\leq n\leq m)$, and using the Lemma 3.4, 3.5, 4.1, 4.6 and the Theorem 4.2, we can obtain the result of this lemma as follows
 $\tau\sum_{n=0}^{m}\|e_{p}^{n+1}\|^{2}\leq C\left(\tau^{2}+h^{2r}\right).$ (4.66)
@@
 Table 2: $L^{2}$- and $H^{1}$-norm errors and convergence orders for velocity and time convergence orders for CHNS (presented by spatial orders).
-$h$ $\left\|\bm{u}-\bm{u}_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\left\|\nabla(\bm{u}-\bm{u}_{h})\right\|_{\ell^{2}(L^{2})}$
+$h$ $\left\|\mathbf{u}-\mathbf{u}_{h}\right\|_{\ell^{\infty}(L^{2})}$ $\left\|\nabla(\mathbf{u}-\mathbf{u}_{h})\right\|_{\ell^{2}(L^{2})}$
 error rate error rate
-1/4  3.4875e--04 -  2.1176e--05  -
-1/8  7.0458e--05 2.3079  3.9048e--06  2.4395
-1/16  8.6735e--06 3.0217  4.9643e--07  2.9754
-1/32  1.0212e--06 3.0861  5.8912e--08  3.0747
-1/64  1.2489e--07 3.0315  7.2101e--09  3.0305
+1/4  3.4875e--04 -  1.6371e--02  -
+1/8  7.0458e--05 2.3079  5.1257e--03  1.6751
+1/16  8.6735e--06 3.0217  1.2321e--03  2.0565
+1/32  1.0212e--06 3.0861  2.9154e--04  2.0794
+1/64  1.2489e--07 3.0315  7.1441e--05  2.0289
 5.2 Coarsening dynamics
@@
 In the experiment, we take $\Omega=[0,1]^{2}$, $\epsilon=0.02$, $M=0.025$, $\nu=0.1$, $\lambda=0.001$, boundary condition on the top is $g=[1,0]$, initial condition for velocity is $\mathbf{u}_{0}=[1,0]$, $\tau=0.001,h=1/64$.
-We solve the problem using scheme (3.6)-(3.12) with the $P_{2}\times P_{2}\times P_{2}\times P_{1}$ for $\phi_{h}^{n+1}$, $\mu_{h}^{n+1}$, $\mathbf{u}_{h}^{n+1}$ and $p_{h}^{n+1}$, snapshots time $t=0,5,10,20,30,50$.
+We solve the problem using scheme (3.6)-(3.12) with the $P_{2}\times P_{2}\times\mathbf{P}_{2}\times P_{1}$ for $\phi_{h}^{n+1}$, $\mu_{h}^{n+1}$, $\mathbf{u}_{h}^{n+1}$ and $p_{h}^{n+1}$, snapshots time $t=0,5,10,20,30,50$.
 The evolution of the concentration for this problem is shown through Figure 4 (see [44] for a similar problem).
@@
 This research is supported by the National Natural Science Foundation of China (No.11971337).
+Appendix
+6.1 Proof of Lemma 3.2
+Proof.
+Firstly, we define a map $\mathcal{G}:S_{h}^{r}\times S_{h}^{r}\rightarrow S_{h}^{r}\times S_{h}^{r}$ by $\mathcal{G}(\phi_{h}^{n+1},\mu_{h}^{n+1})=\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)$, $\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)\in S_{h}^{r}\times S_{h}^{r}$ satisfies
+$\left(\frac{\hat{\phi}_{h}^{n+1}-\phi_{h}^{n}}{\tau},w_{h}\right)-\left(\phi_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla w_{h}\right)+\left(\nabla\hat{\mu}_{h}^{n+1},\nabla w_{h}\right)$ $=0,$ (6.1a)
+$\left(\hat{\mu}_{h}^{n+1},\varphi_{h}\right)-\left(\nabla\hat{\phi}_{h}^{n+1},\nabla\varphi_{h}\right)-\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\varphi_{h}\right)$ $=0,$ (6.1b)
+where $\left(\phi_{h}^{n+1},\mu_{h}^{n+1}\right)\in S_{h}^{r}\times S_{h}^{r}$ and $\left(w_{h},\varphi_{h}\right)\in S_{h}^{r}\times S_{h}^{r}$.
+We will demonstrate that the mapping $\mathcal{G}$ fulfills the conditions of Lemma 3.1, and consequently, it possesses a fixed point that serves as a solution to the scheme (3.6)-(3.7).
+Given $\mathbf{u}_{h}^{n}\in X_{h}^{r+1}$ and $\phi_{h}^{n},\phi_{h}^{n+1}\in S_{h}^{r}$, the CH equation (6.1) can be rewritten as the following problem: find $\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)\in S_{h}^{r}\times S_{h}^{r}$ satisfly
+$a(\hat{\mu}_{h}^{n+1},\varphi_{h})+b(\hat{\phi}_{h}^{n+1},\varphi_{h})=$ $~{}\langle f,\varphi_{h}\rangle,$ (6.2)
+$b(\hat{\mu}_{h}^{n+1},w_{h})-c(\hat{\phi}_{h}^{n+1},w_{h})=$ $~{}\langle g,w_{h}\rangle,$
+for any $\left(w_{h},\varphi_{h}\right)\in S_{h}^{r}\times S_{h}^{r}$, where
+$a(\hat{\mu}_{h}^{n+1},\varphi_{h})=\left(\hat{\mu}_{h}^{n+1},\varphi_{h}\right),~{}b(\hat{\phi}_{h}^{n+1},\varphi_{h})=-\left(\nabla\hat{\phi}_{h}^{n+1},\nabla\varphi_{h}\right),$
+$c(\hat{\phi}_{h}^{n+1},w_{h})=\frac{1}{\tau}\left(\hat{\phi}_{h}^{n+1},w_{h}\right),~{}\langle f,\varphi_{h}\rangle=\left((\phi_{h}^{n+1})^{3}-\phi_{h}^{n},\varphi_{h}\right),$
+$\langle g,w_{h}\rangle=-\frac{1}{\tau}\left(\phi_{h}^{n},w_{h}\right)-\left(\phi_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla w_{h}\right).$
+We denote the $S_{0h}^{r}=\{\varphi_{h}\in S_{h}^{r}:b(\varphi_{h},w_{h})=0,\forall w_{h}\in S_{h}^{r}\}$.
+For any $\varphi_{h}\in S_{0h}^{r}$, it holds $\|\nabla\varphi_{h}\|=0$ and $a(\cdot,\cdot)$ is coercive on $S_{0h}^{r}$.
+Moreover, we can easily show that $a(\cdot,\cdot)$ and $b(\cdot,\cdot)$ are continuous, and $c(\cdot,\cdot)$ is continuous, symmetric and positive semi-definite.
+Thus, according to the Section II.1.2 of [45], given $\mathbf{u}_{h}^{n}\in\mathbf{X}_{h}^{r+1}$ and $\phi_{h}^{n},\phi_{h}^{n+1}\in S_{h}^{r}$, the problem (6.1) is well-posed.
+Furthermore, because the spaces $S_{h}^{r}\times S_{h}^{r}$ are finite dimensional sapces, it holds that $\mathcal{G}$ is a compact map.
+Next, we give the boundedness of $\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)$ in $S_{h}^{r}\times S_{h}^{r}$ as follows:
+$\|\hat{\phi}_{h}^{n+1}\|_{H^{1}}+\|\hat{\mu}_{h}^{n+1}\|_{H^{1}}\leq C,$ (6.3)
+for some $C$ that is a positive constant independent of $\alpha$ and $\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)$.
+Letting $\mathcal{G}(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1})=\frac{1}{\alpha}(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1})$, we have
+$\left(\frac{\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}}{\tau},w_{h}\right)-\alpha\left(\hat{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla w_{h}\right)+\left(\nabla\hat{\mu}_{h}^{n+1},\nabla w_{h}\right)$ $~{}=0,$ (6.4a)
+$\left(\hat{\mu}_{h}^{n+1},\varphi_{h}\right)-\left(\nabla\hat{\phi}_{h}^{n+1},\nabla\varphi_{h}\right)-\alpha\left((\hat{\phi}_{h}^{n+1})^{3}-\phi_{h}^{n},\varphi_{h}\right)$ $~{}=0,$ (6.4b)
+for any $\left(w_{h},\varphi_{h}\right)\in S_{h}^{r}\times S_{h}^{r}$.
+Taking $\left(w_{h},\varphi_{h}\right)=2\left(\tau\hat{\mu}_{h}^{n+1},-\hat{\phi}_{h}^{n+1}+\alpha\phi_{h}^{n}\right)$ into (6.4), and summing the obtained equalities, we get
+$2\tau\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}+\left(\|\nabla\hat{\phi}_{h}^{n+1}\|^{2}-\alpha^{2}\|\nabla\phi_{h}^{n}\|^{2}+\|\nabla(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n})\|^{2}\right)$ (6.5)
+$\frac{\alpha}{2}\left(\|(\hat{\phi}_{h}^{n})^{2}-1\|^{2}-\|(\alpha\phi_{h}^{n})^{2}-1\|^{2}+\|(\hat{\phi}_{h}^{n+1})^{2}-(\alpha\phi_{h}^{n})^{2}\|\right)$
+$\alpha\left(\|\hat{\phi}_{h}^{n+1}(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n})\|^{2}+\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}\right)$
+$\leq$ $~{}2\tau\left(\hat{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla\hat{\mu}_{h}^{n+1}\right)+2\alpha(1-\alpha)\left(\phi_{h}^{n},\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\right)$
+$\leq$ $~{}C\tau\|\hat{\phi}_{h}^{n+1}\|\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}\|\nabla\hat{\mu}_{h}^{n+1}\|+\frac{\alpha}{4}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}+4\alpha(1-\alpha)^{2}\|\phi_{h}^{n}\|^{2}$
+$\leq$ $~{}C\tau\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}+\alpha\phi_{h}^{n}\|^{2}+\tau\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}$
+$~{}+\frac{\alpha}{4}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}+4\alpha(1-\alpha)^{2}\|\phi_{h}^{n}\|^{2}$
+$\leq$ $~{}\frac{\alpha}{4}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}+\frac{\alpha^{2}}{2}\|\phi_{h}^{n}\|^{2}+\tau\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}$
+$~{}+\frac{\alpha}{4}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}+4\alpha(1-\alpha)^{2}\|\phi_{h}^{n}\|^{2}.$
+where we used the following assumption
+$\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}.$ (6.6)
+Remark 6.1.
+According to the assumption $\tau\leq\frac{\alpha}{C\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, the above inequality (6.5) holds.
+Therefore, we can derive the boundedness of the nth-time-layer velocity as follows:
+$\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}\leq C\|\mathbf{u}\|_{L^{\infty}(0,T;\mathbf{L}^{\infty}(\Omega))}\leq C,$ (6.7)
+where $C$ is a positive constant independent of $\alpha$, $\tau$, $h$, and $n$. In this paper, we will analyze the reasonableness of this condition (6.6).
+By from (6.5) and $\alpha\in[0,1]$, we have
+$~{}\tau\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}+\|\nabla\hat{\phi}_{h}^{n+1}\|^{2}+\|\nabla(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n})\|^{2}+\frac{\alpha}{2}\|(\hat{\phi}_{h}^{n+1})^{2}-1\|^{2}$ (6.8)
+$+$ $~{}\alpha\left(\frac{1}{2}\|(\hat{\phi}_{h}^{n+1})^{2}-(\alpha\phi_{h}^{n})^{2}\|^{2}+\|\hat{\phi}_{h}^{n+1}(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n})\|^{2}+\frac{1}{2}\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}\right)$
+$\leq$ $~{}2\|\phi_{h}^{n}\|^{2}+\|\nabla\phi_{h}^{n}\|^{2}+\|(\phi_{h}^{n})^{2}-1\|^{2}\leq C,$
+where $C$ is a positive contant independent of $\left(\hat{\phi}_{h}^{n+1},\hat{\mu}_{h}^{n+1}\right)$ and $\alpha$.
+Setting $\left(w_{h},\varphi_{h}\right)=\left(2\tau\hat{\phi}_{h}^{n+1},2\tau\hat{\mu}_{h}^{n+1}\right)$ in (6.4) and adding the obtained equalities, we get
+$\|\hat{\phi}_{h}^{n+1}\|^{2}-\alpha^{2}\|\phi_{h}^{n}\|^{2}+\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}+2\tau\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}$ (6.9)
+$=$ $~{}2\tau\alpha\left(\hat{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla\hat{\phi}_{h}^{n+1}\right)+2\alpha\tau\left((\hat{\phi}_{h}^{n+1})^{3}-\phi_{h}^{n},\hat{\mu}_{h}^{n+1}\right).$
+Thus, according to the following inequality
+$(\hat{\phi}_{h}^{n+1})^{4}=\left((\hat{\phi}_{h}^{n+1})^{2}-1\right)^{2}+2\left(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\right)^{2}+4\alpha\left(\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\right)\phi_{h}^{n}-1+2\alpha^{2}(\phi_{h}^{n})^{2},$
+we have
+$2\tau\alpha\left(\hat{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla\hat{\phi}_{h}^{n+1}\right)\leq 2\tau\alpha\|\hat{\phi}_{h}^{n+1}\|_{L^{4}}\|\mathbf{u}_{h}^{n}\|_{L^{4}}\|\nabla\hat{\phi}_{h}^{n+1}\|,$ (6.10)
+and
+$2\alpha\tau\left((\hat{\phi}_{h}^{n+1})^{3}-\phi_{h}^{n},\hat{\mu}_{h}^{n+1}\right)\leq 2\tau\alpha\left(\|\hat{\phi}_{h}^{n+1}\|_{L^{4}}^{3}\|\hat{\mu}_{h}^{n+1}\|_{L^{4}}+\|\phi_{h}^{n}\|\|\hat{\mu}_{h}^{n+1}\|\right)$ (6.11)
+$\leq$ $~{}\tau\left(\|\hat{\mu}_{h}^{n+1}\|^{2}+\|\nabla\hat{\mu}_{h}^{n+1}\|^{2}\right)+C_{0}\alpha\tau\left(\|(\hat{\phi}_{h}^{n+1})^{2}-1\|^{2}+\|\hat{\phi}_{h}^{n+1}-\alpha\phi_{h}^{n}\|^{2}\right)^{\frac{3}{2}}$
+$~{}+C_{0}\tau\alpha\left(\alpha^{3}\|\phi_{h}^{n}\|^{3}+\|\phi_{h}^{n}\|^{2}+C\right).$
+Combining the (6.10),(6.11) with (6.9), we deduce the (6.3). ∎
+6.2 Proof of Lemma 3.3
+Proof.
+To prove the existence of a unique solution for the fully discrete scheme (3.6)-(3.12), we suppose the $\left(\phi_{h1}^{n+1},\mu_{h1}^{n+1}\right)$ and $\left(\phi_{h2}^{n+1},\mu_{h2}^{n+1}\right)$ are two solutions of the fully discrete scheme (3.6)-(3.7), and denote
+$\bar{\phi}_{h}^{n+1}=\phi_{h1}^{n+1}-\phi_{h2}^{n+1},\bar{\mu}_{h}^{n+1}=\mu_{h1}^{n+1}-\mu_{h2}^{n+1}.$
+By according to (3.6)-(3.7), we obtain
+$\left(\frac{\bar{\phi}_{h}^{n+1}}{\tau},w_{h}\right)-\left(\bar{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla w_{h}\right)+\left(\nabla\bar{\mu}_{h}^{n+1},\nabla w_{h}\right)$ $~{}=0,$ (6.12a)
+$\left(\bar{\mu}_{h}^{n+1},\varphi_{h}\right)-\left(\nabla\bar{\phi}_{h}^{n+1},\nabla\varphi_{h}\right)-\left(\bar{\phi}_{h}^{n+1}\left((\phi_{h1}^{n+1})^{2}+\phi_{h1}^{n+1}\phi_{h2}^{n+1}+(\phi_{h2}^{n+1})^{2}\right),\varphi_{h}\right)$ $~{}=0.$ (6.12b)
+Letting $\left(w_{h},\varphi_{h}\right)=\left(\tau\bar{\mu}_{h}^{n+1},\bar{\phi}_{h}^{n+1}\right)$ in (6.12) and combining the obtained results, we have
+$\tau\|\nabla\bar{\mu}_{h}^{n+1}\|^{2}+$ $~{}\|\nabla\bar{\phi}_{h}^{n+1}\|^{2}+\left((\phi_{h1}^{n+1})^{2}+\phi_{h1}^{n+1}\phi_{h2}^{n+1}+(\phi_{h2}^{n+1})^{2},(\bar{\phi}_{h}^{n+1})^{2}\right)$ (6.13)
+$=$ $~{}\tau\left(\bar{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla\bar{\mu}_{h}^{n+1}\right).$
+It’s easy to see that
+$\left((\phi_{h1}^{n+1})^{2}+\phi_{h1}^{n+1}\phi_{h2}^{n+1}+(\phi_{h2}^{n+1})^{2},(\bar{\phi}_{h}^{n+1})^{2}\right)$ (6.14)
+$\qquad=\left(\frac{(\phi_{h1}^{n+1})^{2}+(\phi_{h2}^{n+1})^{2}}{2}+\frac{(\phi_{h1}^{n+1}+\phi_{h2}^{n+1})^{2}}{2},(\bar{\phi}_{h}^{n+1})^{2}\right)\geq 0.$
+According to $\tau\leq\frac{C}{\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}}$, we obtain
+$\tau\left(\bar{\phi}_{h}^{n+1}\mathbf{u}_{h}^{n},\nabla\bar{\mu}_{h}^{n+1}\right)\leq$ $~{}\|\bar{\phi}_{h}^{n+1}\|\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}\|\nabla\bar{\mu}_{h}^{n+1}\|$ (6.15)
+$\leq$ $~{}C\tau\|\mathbf{u}_{h}^{n}\|_{L^{\infty}}^{2}\|\bar{\phi}_{h}^{n+1}\|^{2}+\frac{\tau}{2}\|\nabla\bar{\mu}_{h}^{n+1}\|^{2}$
+$\leq$ $~{}\frac{1}{2}\|\nabla\bar{\phi}_{h}^{n+1}\|^{2}+\frac{\tau}{2}\|\nabla\bar{\mu}_{h}^{n+1}\|^{2}.$
+It follows from (6.13) that
+$\|\nabla\bar{\mu}_{h}^{n+1}\|^{2}+\|\nabla\bar{\phi}_{h}^{n+1}\|^{2}=0.$ (6.16)
+Therefore, the discrete scheme (3.6)-(3.8) has a unique solution.
+The uniqueness of the solutions $\left(\mathbf{u}_{h}^{n+1},p_{h}^{n+1}\right)$ for the discrete scheme (3.9)-(3.12) is a standard result.
+Hence, the existence and uniqueness of the solutions for the entire fully discrete scheme (3.6)-(3.12) are proved.
+∎